### v3

# Optimal and Fair Encouragement Policy Evaluation and Learning

Abstract
In consequential domains, it is often impossible to compel individuals to take treatment, so that optimal policy rules are merely suggestions in the presence of human non-adherence to treatment recommendations. Under heterogeneity, covariates may predict take-up of treatment and final outcome, but differently. While optimal treatment rules optimize causal outcomes across the population, access parity constraints or other fairness considerations on who receives treatment can be important. For example, in social services, a persistent puzzle is the gap in take-up of beneficial services among those who may benefit from them the most. We study causal identification and robust estimation of optimal treatment rules, including under potential violations of positivity. We consider fairness constraints such as demographic parity in treatment take-up, and other constraints, via constrained optimization. Our framework can be extended to handle algorithmic recommendations under an often-reasonable covariate-conditional exclusion restriction, using our robustness checks for lack of positivity in the recommendation. We develop a two-stage algorithm for solving over parametrized policy classes under general constraints to obtain variance-sensitive regret bounds. We illustrate the methods in three case studies based on data from reminders of SNAP benefits recertification, randomized encouragement to enroll in insurance, and from pretrial supervised release with electronic monitoring. While the specific remedy to inequities in algorithmic allocation is context-specific, it requires studying both take-up of decisions and downstream outcomes of them.
1 Introduction
Combining causal inference and machine learning to estimate heterogeneous treatment effects can increase revenue and improve public and individual health outcomes by facilitating personalized treatments such as medication decisions or interactions with e-commerce platforms and targeting of social interventions to those who benefit from them most. In particular, a large literature on so-called optimal treatment regimes, or off-policy learning, studies how to leverage treatment effect heterogeneity to identify optimal personalized decision rules for prescriptive decisions about who ought to receive which treatment (Athey, 2017; Kitagawa and Tetenov, 2015; Manski, 2005; Zhao et al., 2012).
However, in many important settings, we cannot compel individuals into treatment. A so-called encouragement design is used to randomize over encouragements, or recommendations into treatment, since it is impossible to randomize the treatment itself. Randomizing treatment may be not possible either because it is ultimately a human-in-the-loop who receives the treatment and it is impossible to compel someone into treatment or because it would be unethical to withhold access to the treatment intervention from the control group. For example, in e-commerce, although companies can assign different users to different visual interfaces of a website, they cannot compel users to sign up for certain services: Instead, they can nudge or encourage users to sign up via promotions and offers. Alternatively, companies may not want to withhold access to a new feature to construct a control group and may instead randomize a new feature’s visibility or the extent of advertising for it (Spotify Engineering, 2023). Then, the treatment group receives more encouragement to use the intervention, while the control group has access but does not receive additional encouragement. Encouragement designs are pervasive not only in e-commerce but more generally in impact assessments across social settings, such as healthcare and social policy. Guidelines for the Abdul Latif Jameel Poverty Action Lab (J-PAL), a major provider of randomized controlled trials in development, note that “when evaluating a noncompulsory, entitlement program, researchers and practitioners cannot and should not force individuals to take up the program nor deny eligible individuals access to the program" and recommend encouragement designs (Heard et al., 2017).
The wide prevalence of encouragement designs implies that optimal treatment regimes also typically operate in the space of encouragements rather than treatment interventions themselves. In particular, we focus on the implications of personalized encouragement interventions for decision-making and fairness or equity concerns; we term such prospective decision rules optimal encouragement policies. We discuss methodological connections with and differences from the well-studied problem of nonadherence in causal inference later on. Often, for the purposes of decision-making, typical approaches adopt an intention-to-treat (ITT) perspective and take the encouragement itself as the treatment (i.e., directly optimizing downstream outcomes of encouragements). For example, pure pricing and profit maximization maximize the ITT effect (i.e., expected profit).
Optimal encouragement policies that personalize encouragement based on covariates are algorithmic allocations that ultimately induce treatment and outcomes. Since we learn about these allocations from historical data, they may suffer issues of algorithmic bias (Barocas et al., 2018). We generally take a multiobjective approach and consider tradeoffs between efficacy (outcomes induced by an optimal encouragement policy) and other equity measures, such as disparities in access (treatment takeup) or encouragement rates across protected groups. ITT-based analysis hides these multiobjective tradeoffs. For example, because of a long history of civil rights oversight, equitable access is of particular concern in public-sector allocations.
Methodologically, the key identifying assumption in evaluations of encouragement designs is what we call the covariate-conditional exclusion restriction (informally, the requirement that the encouragement impact outcomes through no channel other than by increasing treatment takeup), which specifically characterizes this type of design in our view. We give three motivating cross-sector examples that highlight the design’s distinguishing features, namely: 1) Firms can in practice act only in the space of encouragements, not that of treatments themselves; 2) encouragements affect outcomes only by modifying treatment probabilities; 3) firms might have fairness/equity concerns that require multiobjective comparisons; and 4) firms have the operational capacity to potentially target/personalize future encouragements under realistic resource constraints.
Example 1 (Pricing: Demand vs. revenue vs. long-term outcomes).
Firms may consider price discounts for the purposes of long-term outcomes via lead generation or customer acquisition. Price discounts could impact not only short-term revenue but also overall engagement and customer lifetime value. (Karlan and Zinman, 2008) study the impact of price discounts offered by a large microcredit lender. They randomize prices and find that lower interest rates would marginally decrease profits, but they also find evidence of heterogeneity: Women and lower-income individuals take up loans more under discounted rates, with no increase in default. For microcredit lenders interested in financial inclusion, personalized price discounts could expand access while minimizing profit reductions.
Example 2 (Healthcare: Adherence vs. treatment efficacy).
Nonadherence is a major issue in medication treatments and can be significantly affected by social determinants of health.
For example, Miao et al. (2024) studies reasons why patients might switch contraceptives (i.e., move off an initial prescription) and finds significantly higher switching rates for Black and Latinx groups because of not only medical concerns about side effects but also insurance coverage.
Our next example and case study introduces public-sector allocation questions regarding administrative burden in operations, which we revisit in our case study, so we provide additional context and go into more detail. The term administrative burden refers to costs associated with applying for, receiving, and participating in government benefits and services (OMB, 2022). There has been great policy and public-sector interest in reducing administrative burdens. These operational frictions can disproportionately disenfranchise marginalized and disadvantaged groups and can undermine equity-related policy goals. We model several different strategies to reduce administrative burden—reductions in information costs, improvements to outreach via omnichannel and digital marketing, and digital service design (OMB, 2022)—generally through encouragement into treatment (takeup of services).
Example 3 (Takeup of social services and digital outreach).
Many who are eligible for the Supplemental Nutrition Assistance Program (SNAP, formerly known as food stamps) do not receive benefits because of cumbersome processes and understaffed processing centers. While early economic models (Nichols and Zeckhauser, 1982) suggest that those who forgo SNAP application do so because they face a higher time opportunity cost of complying with administrative hassle and thus are likely to be ineligible or to qualify for only small amounts, empirically, this is not the case: Many who are denied benefits because of operational frictions (missed interviews, application errors) are actually eligible. “In 2019, one third of all applications in Los Angeles County were denied due to a missed interview," while five times as many were denied for ineligibility. (Finkelstein and Notowidigdo, 2019) examine behavioral frictions in SNAP benefit access procedures, randomizing both information about likely eligibility and higher-touch outreach via access to a counselor. Though they find overall positive average treatment effects, they also find that "the individuals who apply … in response to either intervention receive lower benefits and are less sick than the average enrollee in the control group”—that is, the applicants for whom the treatment had the intended effect were better off to begin with. How can we instead ensure that interventions reach the worst off, who might benefit most? Under resource constraints, tailoring who receives basic vs. more extensive and expensive outreach can help balance equity with efficiency. Koenecke et al. (2023) find that an efficient advertising budget for GetCalfresh, which streamlines enrollment in California’s CalFresh (SNAP), resulted in low takeup among Hispanic individuals: The authors conduct a survey and find general public support for equitable advertising budget allocations with higher spending on advertising aimed at harder-to-reach populations.
Later, we revisit data from Homonoff and Somerville’s (2021) study of text message reminders about recertification, i.e., the annual benefit renewal procedure. By leveraging heterogeneous effect estimation, we find that the intervention’s efficiency impact is broadly aligned with equity objectives, in the sense that the recipients for whom the treatment effects are greater in terms of next-year benefit dollars are positively rank-correlated with those more likely to respond to the text message reminders. However, personalization should be implemented with caution since this correlation is weaker for nonwhite than for white recipients, giving rise to large access gaps at the top of the distribution.
We develop general methodology that models this structure under a set of causal assumptions, described in full detail later on. Thus far, we have focused on modeling physically randomized encouragements such as those we have just described. However, our model can be extended to model the impacts of algorithmic advice on humans-in-the-loop. In this algorithmic advice setting, the covariate-conditional exclusion restriction is still plausible, but the “overlap" assumption is violated: Randomized (or as-if randomized) algorithmic encouragement or nudge recommendations are typically fixed functions of covariates.
Example 4 (Algorithmic advice).
Doctors prescribe treatment on the basis of algorithmic recommendations (Lin et al., 2021), managers and workers combine their expertise to act on the basis of algorithmic decision support (Bastani et al., 2021), and in the social sector, caseworkers assign individuals to benefit programs on the basis of recommendations from risk scores that support triage (De-Arteaga et al., 2020; Green and Chen, 2019; Yacoby et al., 2022).
Crucially, our model specifically disentangles the effects of algorithmic recommendations on final treatment decisions from the effects of the treatment on outcomes, a contribution that can help make recent approaches based on robustness less conservative. If, as in the settings that we consider here, human decision-makers have wide discretion resulting in treatment overlap, then rather than extrapolating the entire treatment effect function itself, we need extrapolate only the treatment response to the recommendation, which can be more plausible.
Our contributions are as follows. We characterize optimal and resource fairness–constrained optimal decision rules and develop statistically improved estimators and robustness checks for the setting of algorithmic recommendations with sufficiently randomized decisions. In contrast, previous work on algorithmic accountability focuses primarily on auditing recommendations rather than on both the access and efficacy achieved under the final decision rule. Therefore, previous methods can fall short in mitigating potential disparities. We consider two settings: one related to encouragement designs with random allocation of encouragement and another related to algorithmic recommendations (which requires either parametric or robust extrapolation). We also develop methodology for optimizing over a constrained policy class with less conservative out-of-sample fairness constraint satisfaction by means of a two-stage procedure, and we provide sample complexity bounds. We assess the improved recommendation rules in a stylized case study of optimization of health insurance expansion, using data from the Oregon Insurance study, and another stylized case study of optimization of recommendations for pretrial supervised release on the basis of a risk-assessment tool while reducing surveillance disparities.
2 Related Work
In the main text, we briefly highlight the most relevant methodological and substantive work and defer additional discussion to the appendix.
Optimal encouragement designs/policy learning with constraints. There is extensive literature on off-policy evaluation and learning, empirical welfare maximization, and optimal treatment regimes (Athey and Wager, 2021; Zhao et al., 2012; Manski, 2005; Kitagawa and Tetenov, 2015). Qiu et al. (2021) study an optimal individualized encouragement design, though their focus is on optimal individualized treatment regimes with instrumental variables (IVs). Kallus and Zhou (2021a) study fairness in pricing, and some of the desiderata in that setting on revenue (here, marginal welfare) and demand (takeup) are again relevant in our context, but in a more general setting beyond pricing. The results in Kallus and Zhou (2021a) are restricted to pricing and revenue maximization. We consider broader objectives. The most closely related work in terms of problem setup is the formulation of “optimal encouragement designs" in (Qiu et al., 2021). However, they focus on knapsack resource constraints, which have a different solution structure than fairness constraints. Sun et al. (2021) has studied uniform feasibility in constrained resource allocation, but without encouragement or fairness. Ben-Michael et al. (2021) studies robust extrapolation in policy learning from algorithmic recommendation, but not fairness. Our later case study on supervised release leverages the considerable randomness in final treatment decisions for supervised release (decisions that are arguably less consequential than pretrial detention decisions and hence subject to wide discretion) so that we require robust extrapolation over only the first out of two stages.
Fairness constraints in ITT analyses. We focus on deriving estimators for ITT analyses with an eye to relevant fairness constraints. Our interest is in imposing separate desiderata on treatment realizations under noncompliance, but we do not conduct IV inference: We assume that unconfoundedness holds. Our analysis essentially considers simultaneously two perspectives in the constrained optimization: 1) viewing treatment as a potential outcome of a recommendation treatment, i.e., $T(R)$, and 2) taking an ITT stance on the causal effects of treatment on outcomes, i.e., $Y(T)$, even though treatment is not controllable. Marginally, the first perspective yields estimates of disparities and is relevant for estimating fairness constraints, while the second is relevant for estimating the utility objective. Importantly, the quantities that we estimate are not on joint events of takeup and final outcome utility (in contrast to principal stratification). Rather, we assess personalized policies by their population-averaged utility and fairness measures.
3 Problem Setup
We briefly describe the problem setup. We work in the Neyman-Rubin potential outcomes framework for causal inference (Rubin, 1980). We define the following:
  • recommendation flag $R\in\{0,1\}$, where $R=1$ means encouraged/recommended (we use the terms encouragement and recommendation interchangeably);
  • treatment $T(r)\in\mathcal{T}$, where $T(r)=1$ indicates that the treatment decision was $1$ when the recommendation was $r$; and
  • outcome $Y(t(r))$, the potential outcome under encouragement $r$ and treatment $t$.
Regarding fairness, we are concerned about disparities in utility and treatment benefits (resources or burdens) across different groups, denoted $A\in\{a,b\}$. (For notational brevity, we may generically discuss identification/estimation without additionally conditioning on the protected attribute.) For example, recommendations arise from binary high risk/low risk labels of classifiers. In practice, in consequential domains, classifier decisions are rarely automated but rather are used to inform humans-in-the-loop, who decide whether to assign treatment. For binary outcomes, we interpret $Y(t(r))=1$ as the positive outcome. When outcomes and treatments are binary, $Y\in\{0,1\},T\in\mathcal{T}$, where $\mathcal{T}=\{0,1\}$, we may further develop analogues of fair classification criteria. We let $c(r,t,y)\colon\{0,1\}^{3}\mapsto\mathbb{R}$ denote the cost function for $r\in\{0,1\},t\in\mathcal{T},y\in\{0,1\}$, which may sometimes be abbreviated $c_{rt}(y)$. We discuss identification and estimation on the basis of the following recommendation propensity $e_{r}$, treatment propensity $p_{t\mid r}$, and outcome $\mu_{t}$ models:
$e_{r}(X,A)$ $\coloneqq P(R=r\mid X,A),\qquad p_{t\mid r}(X,A)\coloneqq P(T=t\mid R=r,X,A),\;\;$
[…]
We are generally instead interested in personalized recommendation rules, described via the policy function $\pi(r\mid X)\coloneqq\pi_{r}(X)$, which gives the probability of assignment of recommendation $r$ to covariates $X$. The average encouragement effect (AEE) is the difference in average outcomes if we refer everyone vs. no one, while the encouragement policy value $V(\pi)$ is the population expectation induced by the outcomes and treatment with recommendations following the policy distribution.
$AEE=\mathbb{E}[Y(T(1))-Y(T(0))],\qquad V(\pi)=\mathbb{E}[c(\pi,T(\pi),Y(\pi))]$
We use the AEE terminology instead of ITT because the conventional first-stage intention-to-treat in ITT is actually our first-stage encouragement or recommendation. Because algorithmic decision-makers may be differentially responsive to recommendations and treatment effects may be heterogeneous, the optimal recommendation rule may differ from the (infeasible) optimal treatment rule when constraints are taken into account or for simpler policy classes.
Assumption 1 (Consistency and stable unit treatment values [SUTVA).
] $Y_{i}=Y_{i}(T_{i}(R_{i})).$
Assumption 2 (Conditional exclusion restriction).
[…]
Assumption 3 (Unconfoundedness).
[…]
Assumption 4 (Stable responsivities under new recommendations).
$P(T=t\mid R=r,X)$ remains fixed from the observational to the future dataset.
Assumption 5 (Decomposable utilities).
$c(r,t,y)=c_{r}(r)+c_{t}(t)+c_{y}(y)$.
Our key assumption beyond standard causal inference assumptions is assumption 2, the conditional exclusion restriction, i.e., that, conditional on observable information $X$, the recommendation has no causal effect on the outcome beyond its effect on the treatment probability. This assumes that all of the covariate information that is informative of downstream outcomes is measured. Although this assumption may not exactly hold in all applications, stating it is also a starting point for sensitivity analysis under violations of it (Kallus and Zhou, 2018; Kallus et al., 2019b).
Assumption 4 is a structural assumption that limits our method to most appropriately reoptimize over small changes to existing algorithmic recommendations. For example, $p_{0\mid 1}(x)$ (disagreement with algorithmic recommendation) could be a baseline algorithmic aversion. Not all settings are appropriate for this assumption. We do not assume microfoundations on how or why human decision-makers deviate from algorithmic recommendations; rather, we take these patterns as given. Again, we can relax this assumption with sensitivity analysis. Assumption 5 is a mild assumption on modeling utility, namely, that it is not defined on joint realizations of potential outcomes.
We first also assume overlap in recommendations and treatment. Later, however, we give robust methods for relaxing this assumption, leveraging our finer-grained characterization.
Assumption 6 (Overlap).
$\rho_{r}\leq e_{r}(X,A)\leq 1-\rho_{r};\;\;\rho_{t}\leq p_{t\mid r}(X,A)\leq 1-\rho_{t}\text{ and }\rho_{r},\rho_{t}>0$.
4 Method
We consider two problem settings, which model different situations and differ based on the strength of the overlap assumptions.
Setting 1 (Randomized encouragement).
$R$ is (as-if) randomized and satisfies overlap (Assumption 6).
Then, $R$ can be interpreted as the ITT or prescription, whereas $T$ is the actual realization thereof. 1 models nonadherence situations where decision-makers can target encouragements but not direct receipt of the treatment itself.
Setting 2 (Algorithmic recommendation).
$R$ is the output of a predictive model and does not satisfy Assumption 6.
We later extend our methods to the second setting, where $R$ does not satisfy overlap in recommendation but there is sufficient randomness in human decisions to satisfy overlap in treatment.
First, we establish causal identification of the estimands via regression adjustment. Causal identification rewrites the causal estimand in terms of probability distributions estimable from data. The argument follows by our applying the conditional exclusion restriction and consistency but, crucially, does not rely on overlap. We also first consider a special type of fairness constraint, resource parity, and characterize optimal decisions.
Proposition 1 (Regression adjustment identification).
$\mathbb{E}[c(\pi,T(\pi),Y(\pi))]$ $\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mu_{t}(X)p_{t\mid r}(X)]$
Resource parity–constrained optimal decision rules
We consider an access/resource/burden parity fairness constraint:
$V_{\epsilon}^{*}=\min_{\pi}\;\{\mathbb{E}[c(\pi,T(\pi),Y(\pi))]\colon\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]\leq\epsilon\}$ (1)
Enforcing absolute values, etc., follows in the standard way. Not all values of $\epsilon$ may be feasible; in the appendix, we give an auxiliary program for computing feasible ranges of $\epsilon$. We first characterize a threshold solution when the policy class is unconstrained.
Proposition 2 (Threshold solutions).
Define $L(\lambda,X,A)=$
$(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\left\{\tau(X,A)+\frac{\lambda}{p(A)}(\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right])\right\}+\lambda(p_{1\mid 0}(X,a)-p_{1\mid 0}(X,b))$
. Then, $\lambda^{*}\in\arg\min_{\lambda}\;\mathbb{E}[L(\lambda,X,A)_{+}]$ and $\pi^{*}(x,u)=\mathbb{I}\{L(\lambda^{*},X,u)>0\}.$ If instead $d(x)$ is a function of covariates $x$ only, $\lambda^{*}\in\arg\min_{\lambda}\;\mathbb{E}[\mathbb{E}[L(\lambda,X,A)\mid X]_{+}]$ and $\pi^{*}(x)={\mathbb{I}\{\mathbb{E}[L(\lambda^{*},X,A)\mid X]>0\}}.$
Establishing this threshold structure (which follows by duality of infinite-dimensional linear programming) allows us to provide a generalization bound argument.
Proposition 3 (Policy value generalization).
Assume that the nuisance models $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0}]^{\top},\eta\in\mathcal{F}_{\eta}$ are consistent and well specified with finite Vapnik–Chervonenkis (VC) dimension $v_{\eta}$ over the product function class $\mathcal{F}_{\eta}$. Let $\Pi=\{\mathbb{I}\{\mathbb{E}[L(\lambda,X,A;\eta)\mid X]>0\colon\lambda\in\mathbb{R};\eta\in\mathcal{F}_{\eta}\}.$
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi L(\lambda,X,A)]-\mathbb{E}[\pi L(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
This bound is stated for known nuisance functions; verifying stability under estimated nuisance functions further requires rate conditions.
Doubly robust estimation
We may improve the statistical properties of the estimation by developing doubly robust estimators, which can achieve faster statistical convergence when both the probability of recommendation assignment (when it is random) and the probability of outcome are consistently estimated or can otherwise protect against misspecification of either model. We first consider the ideal setting when algorithmic recommendations are randomized so that $e_{r}(X)=P(R=r\mid X)$.
Proposition 4 (Variance-reduced estimation).
$\textstyle V(\pi)$ $=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\frac{\mathbb{I}\left[R=r\right]}{e_{r}(X)}(\mathbb{I}[T=t]c_{rt}(Y)-\mu_{t}(X)p_{t\mid r}(X))+\mu_{t}(X)p_{t\mid r}(X)\right\}\right]$
$\mathbb{E}[T(\pi)]$ $=\textstyle\sum_{r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\frac{\mathbb{I}\left[R=r\right]}{e_{r}(X)}(T(r)-p_{1\mid r}(x))+p_{1\mid r}(x)\right\}\right]$
We retain the full expression rather than simplifying (as appears in Qiu et al., 2021) since the doubly robust estimation of constraints changes the Lagrangian. For example, for regression adjustment, it is clearer in Proposition 9 how constraints affect the optimal decision rule.
4.1 Robust estimation with treatment overlap but without recommendation overlap
When the recommendations are, e.g., the high risk/low risk labels from binary classifiers, the overlap assumption may not be satisfied since the algorithmic recommendations are deterministic functions of covariates. However, note that identification in Proposition 1 requires only SUTVA, consistency, and the exclusion restriction.
A naive approach based on parametric extrapolation is to estimate $p_{1\mid 1}(X)$, treatment responsivity, on the observed data and simply use the parametric form to extrapolate to the full dataset. (In Appendix B, we describe the variance reduction that can be possible). On the other hand, parametric extrapolation is generally unsatisfactory because the conclusions will be driven by the model specification rather than by observed data. Nonetheless, it can provide a starting point for robust extrapolation of structurally plausible treatment response probabilities.
Robust extrapolation under violations of overlap
We next describe methods for robust extrapolation under structural assumptions about the smoothness of the outcome models. Under violations of overlap, the only unknown quantity is $p_{t\mid r}(X)$ in regions of no overlap in recommendation; however, a plausible assumption is that the underlying function is smooth in covariates. A robust approach obtains worst-case bounds on policy value under all functions compatible with a particular smoothness assumption. On the other hand, we assume that overlap holds with respect to $T$ given covariates $X$, so our finer-grained approach via Assumption 2 yields milder penalties due to robustness since we need robustly extrapolate only the treatment response to recommendations, $p_{t\mid r}(X),$ rather than the outcome model’s $\mu_{t}(X).$ Define the regions of no overlap as the following: Let $\mathcal{X}^{\text{nov}}_{r}=\{x:P(R=r\mid x)=0\}$; in this region, we do not jointly observe all potential values of $(t,r,x)$. In addition, let $\mathcal{X}^{\text{nov}}=\bigcup_{r}\mathcal{X}^{\text{nov}}_{r}$. Correspondingly, define the overlap region as $\mathcal{X}^{\text{ov}}=(\mathcal{X}^{\text{nov}})^{c}.$ We consider uncertainty sets for ambiguous treatment recommendation probabilities. For example, one plausible structural assumption is monotonicity of treatment in recommendation. We define the following uncertainty set:
$\mathcal{U}_{q_{t\mid r}}\coloneqq\left\{q_{1\mid r}(x^{\prime})\colon q_{1\mid r}(x)\geq p_{1\mid r}(x),\;\forall x\in\mathcal{X}^{\text{nov}}_{r}\;\sum_{t\in\mathcal{T}}q_{t\mid r}(x)=1,\forall x,r\right\}$
We could assume uniform bounds on unknown probabilities; more refined bounds, such as Lipschitz smoothness with respect to some distance metric $d$; or boundedness.
$\textstyle\mathcal{U}_{\text{lip}}$$\coloneqq\left\{q_{1\mid r}(x^{\prime})\colon d(q_{1\mid r}(x^{\prime}),p_{1\mid r}(x))\leq Ld(x^{\prime},x),\;(x^{\prime},x)\in(\mathcal{X}^{\text{nov}}\times\mathcal{X}^{\text{nov}})\right\}$
$\mathcal{U}_{\text{bnd}}$$\coloneqq\left\{q_{1\mid r}(x^{\prime})\colon\underline{b}(x)\leq q_{1\mid r}(x^{\prime})\leq\overline{b}(x)\right\}$
Define $V_{ov}(\pi)\coloneqq\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)p_{t\mid r}(X)\mu_{t}(X)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}]$. Let $\mathcal{U}$ denote the uncertainty set including any custom constraints, e.g., $\mathcal{U}=\mathcal{U}_{q_{t\mid r}}\cap\mathcal{U}_{\text{lip}}$. Then, we may obtain robust bounds by optimizing over regions of no overlap:
$\textstyle\overline{V}(\pi)$ $\coloneqq V_{ov}(\pi)+\overline{V}_{nov}(\pi),\;\;$
[…]
In the specialized but practically relevant case of binary outcomes/treatments/recommendations, we obtain the following simplifications for bounds on the policy value and the minimax robust policy that optimizes the worst-case overlap extrapolation function. In the special case of constant uniform bounds, it is equivalent (in the case of binary outcomes) to consider marginalizations:
Lemma 1 (Binary outcomes, constant bound).
Let $\mathcal{U}_{\text{cbnd}}\coloneqq\left\{q_{t\mid r}(x^{\prime})\colon\underline{B}\leq q_{1\mid r}(x^{\prime})\leq\overline{B}\right\}$ and $\mathcal{U}=\mathcal{U}_{q_{t\mid r}}\cap\mathcal{U}_{\text{cbnd}}.$ Define $\beta_{t\mid r}\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t].$If $T\in\{0,1\},$
$\overline{V}_{no}(\pi)$ $=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[c^{*}_{rt}\beta_{t\mid r}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]],$
[…]
We consider the case of continuous-valued outcomes in the example setting of the simple resource parity–constrained program of Equation 2. We first study simple uncertainty sets, such as intervals, to deduce insights about the robust policy, with a more general reformulation in the appendix.
Proposition 5 (Robust linear program).
Suppose that $r,t\in\{0,1\}$ and $q_{r1}(\cdot,u)\in\mathcal{U}_{bnd},\forall r,u.$ Define
$\tau(x,a)$ $\coloneqq\mu_{1}(x,a)-\mu_{0}(x,a),\;\;\Delta B_{r}(x,u)\coloneqq(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)),$
$B^{\text{mid}}_{r}(x,u)$ $\coloneqq\underline{B}_{r}(x,u)+\frac{1}{2}\Delta B_{r}(x,u),\;\;c_{1}(\pi)\textstyle\coloneqq\sum_{r}\mathbb{E}[\tau\pi_{r}B^{\text{mid}}],$
$\mathbb{E}[\Delta_{ov}T(\pi)]$ $\coloneqq\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}\mid A=b]\;\;$
Then, the robust linear program is:
$\min\;$ $\textstyle V_{ov}(\pi)+\mathbb{E}[\mu_{0}]+c_{1}(\pi)-\frac{1}{2}\sum_{r}\mathbb{E}[\left|\tau\right|\pi_{r}\Delta B_{r}(X,A)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]$
s.t. $\textstyle\sum_{r}\{\mathbb{E}[\pi_{r}\overline{B}_{r}(X,A)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}\mid A=a]-\mathbb{E}[\pi_{r}\underline{B}_{r}(X,A)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}\mid A=b]\}+\Delta_{ov}^{T}(\pi)\leq\epsilon$
5 Additional Fairness Constraints and Policy Optimization
We previously discussed policy optimization over unrestricted decision rules given estimates. We now introduce a general methodology to handle 1) optimization over a policy class of restricted functional form and 2) more general fairness constraints. We first introduce the fair-classification algorithm of Agarwal et al. (2018) and then describe our extensions to obtain variance-sensitive regret bounds and less conservative policy optimization (using a regularized empirical risk minimization (ERM) argument given in Chernozhukov et al., 2019).
Algorithm and setup
Algorithm 1 $\text{REDFAIR}(\mathcal{D},g,\mathcal{E},M,d)$
1:  Input: $\mathcal{D}=\left\{\left(X_{i},R_{i},T_{i},Y_{i},A_{i}\right)\right\}_{i=1}^{n}$, $g,\mathcal{E},{M},\widehat{{d}}$, B, accuracy $\nu$, $\alpha$, stepsize $\omega,$ initialization ${\theta}_{1}={0}\in\mathbb{R}^{|\mathcal{K}|}$
2:  for iteration $t=1,2,\dots$  do
3:     Set $\lambda_{t,k}=B\frac{\exp\left\{\theta_{k}\right\}}{1+\sum_{k^{\prime}\in\mathcal{X}}\exp\left\{\theta_{k^{\prime}}\right\}}$ for all $k\in\mathcal{K}$, $\beta_{t}\leftarrow\operatorname{BEST}_{\beta}\left({\lambda}_{t}\right)$, $\widehat{Q}_{t}\leftarrow\frac{1}{t}\sum_{t^{\prime}=1}^{t}\beta_{t^{\prime}}$ $\hat{{\lambda}}_{t}\leftarrow\frac{1}{t}\sum_{t^{\prime}=1}^{t}{\lambda}_{t^{\prime}},$
4:     $\bar{L}\leftarrow L(\hat{Q}_{t},\operatorname{BEST}_{{\lambda}}(\hat{Q}_{t})),\underline{L}\leftarrow(\operatorname{BEST}_{\beta}(\hat{{\lambda}}_{t}),\hat{{\lambda}}_{t})$,
5:     $\nu_{t}\leftarrow\max\{L(\hat{Q}_{t},\hat{{\lambda}}_{t})-\underline{L},\quad\bar{L}-L(\hat{Q}_{t},\hat{{\lambda}}_{t})\}$, If $\nu_{t}\leq\nu$ then return $(\hat{Q}_{t},\hat{{\lambda}}_{t})$
6:     $\theta_{t+1,i}=\theta_{t}+\log(1-\omega({M}\hat{{\mu}}\left(h_{t}\right)-\hat{{c}})_{j}),\forall i$
7:  end for
We first describe the reductions-based approach for fair classification of Agarwal et al. (2018) before describing our adaptation for constrained policy learning and localized two-stage variance reduction. They consider classification (i.e., loss minimization) under fairness constraints that can be represented generically as a linear program. In the following, note that, to be consistent with standard form for linear programs, we consider costs $Y$ so that we can phrase the saddle point as minimization-maximization. The $|\mathcal{K}|$ linear constraints and $J$ groups (values of protected attribute $A$) are summarized via a coefficient matrix ${M}\in\mathbb{R}^{K\times J}$, which multiplies a vector of constraint moments $h_{j}(\pi),j\in[J]$ (with $J$ being the number of groups); $O=(X,A,R,T,Y)$ denotes our data observations and $d$ the constraint constant vector:
$h_{j}(\pi)=\mathbb{E}\left[g_{j}(O,\pi(X))\mid\mathcal{E}_{j}\right]\;\;\text{ for }j\in J,\;\;{M}{h}(\pi)\leq{d}$
The elements of $h_{j}(\pi)$ are average functionals (for example, the average treatment takeup in group $j$). Importantly, the moment function $g_{j}$ depends on $\pi$, while the conditioning event $\mathcal{E}_{j}$ cannot depend on $\pi$. Many important fairness constraints can nonetheless be written in this framework, such as burden/resource parity and parity in true positive rates, but not measures such as calibration whose conditioning event does depend on $\pi$. (See Section C.2 for examples omitted for brevity.)
Our objective function is the policy value $V(\pi)$. (Later, this is linearized (as in Agarwal et al., 2018) by optimizing over distributions over policies.) We further consider a convexification of $\Pi$ via randomized policies $Q\in\Delta(\Pi)$, where $\Delta(\Pi)$ is the set of distributions over $\Pi,$ i.e., a randomized classifier that samples a policy $\pi\sim Q$. Therefore, our target estimand is the optimal distribution $Q$ over policies $\pi$ that minimizes the objective value $V(Q)$ subject to the fairness constraints encoded in ${M}{h}(Q)\leq{d}$:
$\min_{Q\in\Delta(\Pi)}\{V(Q)\colon\;\;{M}{h}(Q)\leq{d}\}$
Next, we discuss the cost-weighted classification reduction of off-policy learning (Zhao et al., 2012), which we use to solve constrained off-policy learning via Agarwal et al. (2014).
We use a well-known reduction of policy learning to cost-sensitive classification, described in Section C.2.1 of the appendix. Therefore, the centered regret can be reparametrized via the parameter $\beta$ as: $J(\beta)=J(\operatorname{sgn}(f_{\beta}(\cdot)))=\mathbb{E}[\operatorname{sgn}(f_{\beta}(X))\left\{\psi\right\}].$ We can apply the standard reduction to cost-sensitive classification since $\psi_{i}\operatorname{sgn}(f_{\beta}(X_{i}))=\left|\psi_{i}\right|(1-2\mathbb{I}\left[\operatorname{sgn}(f_{\beta}(X_{i}))\neq\operatorname{sgn}(\psi_{i})\right]$). Then, we can use surrogate losses for the zero-one loss. Although many functional forms for $\ell(\cdot)$ are Fisher consistent, one such choice of $\ell$ is the logistic (cross-entropy) loss $\mathbb{E}[\left|\psi\right|\ell(f_{\beta}(X),\operatorname{sgn}(\psi))],\qquad l(g,s)=2\log(1+\exp(g))-(s+1)$.
Algorithm 2 Two-stage localized fair classification via reductions
1:  Randomly split the data into two folds, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
2:  Obtain $\hat{Q}^{*}_{1}$ and the index set of binding constraints ^$\hat{\mathcal{I}}_{1}$ by learning nuisances $\eta_{1}$ and running Algorithm 1 on $\mathcal{D}_{1}$ with $\text{REDFAIR}(\mathcal{D}_{1},h,{\mathcal{E}},M,d;\eta_{1}).$
3:  $\hat{\sigma}^{2}_{j}\leftarrow\hat{\operatorname{Var}}(g_{j}(O,\hat{Q}_{1})\mathbb{I}\left[\mathcal{E}_{j}\right]/p_{j}),\;\forall j$ $\hat{d}\leftarrow d+2\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\hat{\sigma}^{2}_{j}n^{-\alpha}$
4:  Augment additional constraints with $\epsilon_{n}$ policy value and constraint slices relative to $\hat{\pi}_{1}$; define an augmented system (where subscripting by ^$\hat{\mathcal{I}}_{1}$ subindexes the corresponding matrix or vector):
$\tilde{h}_{j^{\prime}}(Q)$ $=\mathbb{E}_{n_{1}}[\{g_{j^{\prime}}(O;\hat{Q}_{1})-g_{j^{\prime}}(O;Q)\}\mid\mathcal{E}_{j}],\;\;\forall j^{\prime}\in\hat{\mathcal{I}}_{1},\;$
$\tilde{h}^{v}(Q)$ $=\mathbb{E}_{n_{1}}[v_{DR}(O;\hat{Q}_{1},\eta_{1})-v_{DR}(O;Q,\eta_{1})]$
$\tilde{M}$ $=[M;M_{\hat{\mathcal{I}}_{1}},\vec{1}],\tilde{h}=[h,\tilde{h},\tilde{h}^{v}]^{\top},\tilde{d}=[\hat{d},\epsilon_{n}\vec{1},\epsilon_{n}]^{\top},\tilde{\mathcal{E}}=[\mathcal{E},\mathcal{E}_{\hat{\mathcal{I}}_{1}},\emptyset]^{\top}$
5:  Solve $\min_{Q\in\Delta(\Pi)}\{\hat{V}(Q)\colon\;\;\tilde{M}\tilde{h}(Q)\leq\tilde{d}\}.$ Obtain $\hat{Q}^{*}_{2}$ by running Algorithm 1 on $\mathcal{D}_{2}$ with $\text{REDFAIR}(\mathcal{D},\tilde{g},\tilde{\mathcal{E}},\tilde{M},\tilde{d},\eta_{2})$.
Optimization. Ultimately, the optimization is solved with sampled and estimated moments. Define the integrand of the constrained, weighted empirical risk minimization as $v_{(\cdot)}(O;\pi_{\beta},\eta)=\left|\psi_{(\cdot)}(O;\eta)\right|\ell(f_{\beta}(X),\operatorname{sgn}(\psi_{(\cdot)}(O;\eta))).$ Our estimate of the objective function is therefore
$V_{(\cdot)}(Q)=\mathbb{E}[\left|\psi_{(\cdot)}\right|\ell(f_{\beta},\operatorname{sgn}(\psi_{(\cdot)}))]=\mathbb{E}_{\pi_{\beta}\sim Q}[v_{(\cdot)}(O;\pi_{\beta},\eta)].$
Note that for the rest of our discussions of algorithms for constrained policy optimization, we overload notation and use $V_{(\cdot)}(Q)$ to refer to policy regret, as above. The optimal policies are the same for regret and for value. We obtain the sample estimator $\hat{V}_{(\cdot)}(Q)$ and sample constraint moments $\hat{h}(Q)$ analogously. We also add a feasibility margin $\epsilon_{k}$ that depends on concentration of the estimated constraints, so the sampled constraint vector is $\hat{d}_{k}=d_{k}+\epsilon_{k},$ for all $k$. We seek an approximate saddle point so that the constrained solution is equivalent to the Lagrangian,
$\hat{L}(Q,{\lambda})=\hat{V}(Q)+{\lambda}^{\top}({M}\hat{{h}}(Q)-\hat{{d}}),\qquad\min_{Q\in\Delta(\Pi)}\{\hat{V}(Q)\colon M\hat{h}(Q)\leq\hat{d}\}=\min_{Q\in\Delta(\Pi)}\max_{{\lambda}\in\mathbb{R}_{+}^{K}}\hat{L}(Q,\lambda).$
We simultaneously solve for an approximate saddle point over the $B$-bounded domain of $\lambda$:
$\textstyle\underset{Q\in\Delta}{\min}\;\;\underset{{{\lambda}\in\mathbb{R}_{+}^{|\mathcal{X}|},\|{\lambda}\|_{1}\leq B}}{\max}\hat{L}(Q,{\lambda}),\qquad\underset{{\lambda}\in\mathbb{R}_{+}^{|\mathcal{X}|},\|{\lambda}\|_{1}\leq B}{\max}\;\;\underset{Q\in\Delta}{\min}\;\hat{L}(Q,{\lambda})$
(Agarwal et al., 2018, Theorem 3) gives generalization guarantees on the policy value and constraint violation achieved by the approximate saddle point output by the algorithm. The analysis is generic under rate assumptions on uniform convergence of policy and constraint values. Such a rate $\alpha$ follows from standard analyses in causal inference and is used to set the constraint violation feasibility margin $\epsilon_{k}=O(n^{-\alpha})$.
Assumption 7 (Rate assumption on policy and constraint values).
There exist $C,C^{\prime}\geq 0$ and $\alpha\leq 1/2$ such that $\sup_{Q\in\Delta(\Pi)}\{V(Q;\eta)-\hat{V}(Q;\hat{\eta})\}\leq Cn^{-\alpha}$ and $\varepsilon_{k}=C^{\prime}\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|n_{j}^{-\alpha}$, where $n_{j}$ is the number of data points that fall in $\mathcal{E}_{j}$.
Next, we summarize the optimization algorithm. We play a no-regret (second-order multiplicative weights; Cesa-Bianchi et al., 2007; Steinhardt and Liang, 2014) algorithm (a slight variant of the hedge/exponentiated gradient algorithm; Freund and Schapire, 1997) for the $\lambda$-player while using best-response oracles for the $Q$-player. Full details are in Algorithm 1. Given $\lambda_{t},$ $\operatorname{BEST}_{\beta}\left({\lambda}_{t}\right)$ computes a best response over $Q$; since the worst-case distribution will place all its weight on one classifier, this step can be implemented by a reduction to cost-sensitive/weighted classification (Beygelzimer and Langford, 2009; Zhao et al., 2012), which we describe in further detail below. Computing the best response over $\operatorname{BEST}_{{\lambda}}(\hat{Q}_{t}))$ selects the most violated constraint. Further details are in Section C.2.
Two-stage variance-constrained algorithm. We seek to improve upon this procedure so that we may obtain regret bounds on policy value and fairness constraint violation that exhibit more favorable dependence on the maximal variance over small-variance slices near the optimal policy, rather than worst-case constants over all policies (Chernozhukov et al., 2019; Athey and Wager, 2021). Moreover, using the estimated variance to set constraint feasibility slacks can achieve tighter fairness control.
These challenges motivate the two-stage procedure, described formally in Algorithm 2 and verbally here. We adapt an out-of-sample regularization scheme developed in (Chernozhukov et al., 2019), which recovers variance-sensitive regret bounds via a small modification to an ERM procedure (and, by extension, policy learning). We split the data into two subsets, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and first learn nuisance estimators $\hat{\eta}_{1}$ from $\mathcal{D}_{1}$ (possibly with further sample splitting) for use in our policy value and constraint estimates. We run Algorithm 1 ($\operatorname{REDFAIR}(\mathcal{D}_{1},h,{\mathcal{E}},M,d;\hat{\eta}_{1})$) on data from $\mathcal{D}_{1}$ to estimate the optimal policy distribution $\hat{Q}_{1}$ and the constraint variances at $\hat{Q}_{1}$. We identify the first-stage binding constraints via the index set $\hat{I}_{1}$. Next, we augment the constraint matrix with additional constraints that require feasible policies for the second-stage policy distribution to achieve $\epsilon_{n}$ close policy value and constraint moment values relative to $\hat{Q}_{1}$. Since errors concentrate quickly, this can be viewed as variance regularization. In addition, we set the constraint slacks $\hat{d}\leftarrow d+2\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\hat{\sigma}^{2}_{j}n^{-\alpha}$ in the second stage using the estimated variance constants from $\hat{Q}_{1}$. This results in tighter control of the fairness constraints. The second stage solves for an approximate saddle point of the augmented system, with objective function and constraints evaluated on $\mathcal{D}_{2}$ and returns $\hat{Q}_{2}.$
[…]
Assumption 8.
The function classes $\mathcal{F}_{\Pi},\{\mathcal{F}_{j}\}_{j\in\mathcal{J}}$ satisfy that, for any constant $r,\kappa(r,\mathcal{F})\rightarrow 0$ as $n\rightarrow\infty$. The function classes $\{\mathcal{F}_{j}\}_{j\in\mathcal{J}}$ comprise $L_{j}$-Lipschitz contractions of $\pi.$
We assume that we are using doubly robust/orthogonalized estimation as in proposition 4 and, hence, state our results depending on the estimation error of nuisance vector $\eta$. The next theorem summarizes the out-of-sample performance of the two-stage algorithm of Algorithm 2, $\hat{Q}_{2}$.
Theorem 3 (Variance-based oracle policy regret).
Suppose that the mean-squared error of the nuisance estimates is upper bounded w.p. $21-\delta/$ 2 by $\chi_{n,\delta}^{2}$ over the randomness of the nuisance sample: $\max_{l}\{\mathbb{E}[(\hat{\eta}_{l}-\eta_{l})^{2}]\}_{l\in[L]}\coloneqq\chi_{n}^{2}$
Let $v_{DR}^{0}(O;Q)$ denote evaluation with true nuisance functions $\eta_{0}$; define $r=\sup_{Q\in\mathcal{Q}}\sqrt{\mathbb{E}\left[v_{DR}^{0}(O;Q)^{2}\right]}$ and $\epsilon_{n}=\Theta\left(\kappa\left(r,\mathcal{F}_{\Pi}\right)+r\sqrt{\frac{\log(1/\delta)}{n}}\right)$. Moreover, denote an $\epsilon$-regret slice of the policy space:
$\mathcal{Q}_{*}(\epsilon)=\left\{Q\in\Delta[\Pi]:V(Q_{*}^{0})-V(Q)\leq\epsilon,\;\;h(Q_{*}^{0})-h(Q)\leq d+\epsilon\right\}$
Let $\tilde{\epsilon}_{n}=O(\epsilon_{n}+\chi_{n,\delta}^{2})$, and denote the variance of the difference between any two policies in an $\epsilon_{n}$-regret slice, evaluated at the true nuisance quantities:
$\bar{\sigma}_{\mathcal{D}_{2}}^{2}=\sup\;\{\operatorname{Var}\left(v_{DR}^{0}(O;Q)-v_{DR}^{0}\left(O;Q^{\prime}\right)\right)\colon{Q,Q^{\prime}\in\mathcal{Q}_{*}(\tilde{\epsilon}_{n})}\}$
(Define $\bar{\sigma}_{k,\mathcal{D}_{2}}^{2}$ analogously for the variance of constraint moments.) Then, letting ${{\gamma}}(Q):={M}{{h}}(Q)$ denote the constraint values, the policy distribution $Q_{2}$ returned by the out-of-sample regularized ERM satisfies w.p. $1-\delta$ over the randomness of $S$:
$V(\hat{Q}_{2})-V(Q^{*})$ $=\textstyle O(\kappa(\bar{\sigma}_{\mathcal{D}_{2}},\operatorname{conv}(\mathcal{F}_{\Pi}))+\bar{\sigma}_{\mathcal{D}_{2}}{n}^{-\frac{1}{2}}\sqrt{{\log(3/\delta)}}+\chi_{n,\delta}^{2})$
$(\gamma_{k}(\hat{Q}_{2})-d_{k})-(\gamma_{k}(Q^{*})-d_{k})$ $=\textstyle O(\kappa(\bar{\sigma}_{k,\mathcal{D}_{2}},\operatorname{conv}(\mathcal{F}_{j}))+\bar{\sigma}_{k,\mathcal{D}_{2}}{n}^{-\frac{1}{2}}\sqrt{{\log(3/\delta)}}+\chi_{n,\delta}^{2})$
The specific benefits of the two-stage approach are that 1) the constants are improved from being absolute, structure-agnostic bounds to depending on the variance of low-regret policies, which also reflects the improved variance from the use of doubly robust estimation as in proposition 4, and 2) it allows less conservative satisfaction of the fairness constraint out of sample.
6 Connections to IV Estimation
Nonadherence has classically been studied in econometrics with IVs. In this work, we focus primarily on the implications of Assumption 2. Common IV analyses impose additional assumptions such as monotonicity or no defiers; these can impose restrictions in behavior that empirical studies show are sometimes untrue. Nonetheless, in other domains, these assumptions can be more or less plausible, and so, in this section, we discuss connections to other IV models for completeness.
However, a primary point of our paper is that the (potentially covariate-conditional) encouragement effect is the relevant estimand for decision-making purposes. One naive approach to decision-making is the following: Estimate a conditional local average treatment effect with advanced methodology; under monotonicity, compliers drive the encouragement effect, so try to classify who is a complier, and give them the encouragement. An immediate issue here is that complier classification based on the compliance score does not account for the downstream impacts of correct/incorrect classification of compliers and so the classification ought to be weighted by decision-theoretic utility (the clustered local average treatment effect [CLATE]). Making this adjustment just is the conditional encouragement estimand that we focus on in this paper. Further, auditing such a two-stage procedure for fairness introduces complications, as complier classification itself can introduce fair classification concerns. Disaggregating the impacts on takeup/compliance with service enrollment and downstream outcomes can reveal differing sources of inequity and point to different directions for further research.
Fair complier classification auditing under monotonicity
7 Case Studies
We first include an empirical case study that analyzes heterogeneous treatment effects and compliance effects in Homonoff and Somerville (2021), which demonstrates the importance of disaggregated evaluations of efficacy and takeup. Next, we consider approaches based on constrained policy optimization.
7.1 Case study: Text message reminders
We consider a case study in the context of text message reminders of recertification interviews for SNAP, a benefit program providing electronic transfers for food purchases, formerly known as food stamps. We first discuss background on the data and the context of the text message reminders, which are expected to increase compliance with the benefit recertification process. Next, we describe our focal analysis task: reoptimizing reminders to optimize impacts (encouragement) on attendance of a recertification interview (treatment) and therefore on next-year benefits (outcome). We first provide a descriptive analysis outlining the heterogeneous effects of encouragement on treatment and of treatment on outcomes, with a focus on comparing group differences. Next, we turn to off-policy evaluation and optimization, considering budget allocations and where potential equity concerns may arise in the distribution of final outcomes ($\mathbb{E}[Y(\pi)\mid A=a]$) vs. the distribution of improvements.
7.1.1 Background
Applications for public services are complex. Federal guidelines require an interview, during which a caseworker typically goes over the application and explains to the applicant complex eligibility guidelines and rules, such as idiosyncratic income reporting and household determination guidelines. The interview can be completed in person or over the phone, after which individuals must provide requested documents within ten days. Recent digital innovations (such as GetCalFresh) have improved the application process. Exact operational delivery is the responsibility of state and local administrators, with some variation in procedural features such as flexible callback systems, scheduling of interviews, or central management of call-in interviews by a call center. We focus on the recertification interview: Every year, to renew benefits, households must conduct another interview to recertify their eligibility, in case of changes to their income or household composition. In San Francisco, California, the case study location, interviews were scheduled completely at random over the course of the next month and could be conducted in person or by phone. In-person interviews impose a hassle cost on the applicant, for example, due to transportation or childcare costs, inflexible work schedules, or other precarities experienced by low-income households applying for benefits.
We reanalyze the data from Homonoff and Somerville (2021), focusing on the text message reminders. Homonoff and Somerville (2021) are primarily interested in the effect of the time between the notice of the scheduled interview and the interview itself on individuals’ recertification and downstream outcomes, such as yearly SNAP benefits. Overall, they find positive impacts of the time lapse between the scheduling of the interview and the end of the month when the recertification process should be completed. The data also record the results from a contemporaneous pilot study of text message reminders about the recertification deadline, which is our focus. Individuals could opt into a text message reminder that their recertification was due at the end of the month and directing them to call the San Francisco Human Services Agency (SFHSA) in case of questions. In addition to the interview, recertification requires the individuals gather certain documents to prove eligibility, with intermediate deadlines for these other processes. The reminder about the overall deadline is theorized to improve awareness of time-sensitive aspects of recertification and improve interview compliance. The original paper (Homonoff and Somerville, 2021) finds a positive average treatment effect. Important questions remain, however, about potential heterogeneity in the impacts of the text message reminders on interview completion (and of interview completion on yearlong outcomes). We focus on the question of reoptimizing textmessage nudges. We leave the question of joint optimization of text message reminders and interview time for future work and assume that current scheduling operations remain fixed. The interview date within the next month is assigned completely at random to balance workload. Assuming that individuals arrive randomly throughout the month, it is plausible that the distribution of the time lapse prior to interview date itself remains random.
Text message reminders are a low-touch, tactical lever to reduce behavioral frictions and costs that could lead to missed interviews. If a beneficiary fails to complete the interview within the given 30-day time window, her benefits are terminated, and although households can later reschedule, benefits are not prorated, leading to benefit losses. In our analysis, we focus on potential optimization of text message reminders and equity concerns. Given that text message reminders are unlikely to reduce interview attendance rates, without resource constraints, the results of Homonoff and Somerville (2021) already suggest they should be made available to all. On the other hand, resource constraints are often binding for nonprofits and government, and thus cost optimization can be critical: Eligibility determinations affect wide swathes of the population, and more intensive outreach options could further strain budgets. For example, the Benefits Data Trust (BDT), a longtime service provider, recently declared the end of its operational model, a move speculated to be due in part to high operational costs.
Our methods illustrate the tradeoffs characterizing different potential targeting rules. For example, our methodology allows more precise estimates of the average encouragement effect, which can be used to evaluate the value of the so-called default option question in nudge designs: What are the gains from a move from opting in to opting out of text message reminders? Could the global budget constraints necessary in practice inadvertently introduce inequality?
7.1.2 Model and descriptive statistics: Heterogeneous takeup and efficacy
Next, we describe our specific instantiations of our general encouragement model. We let $R\in\{0,1\}$ denote receipt of a text message reminder, where $R=1$ denotes that a beneficiary received a reminder. We consider the treatment $T\in\{0,1\}$, where $T=1$ indicates attendance of any recertification interview (enrollees can reschedule another interview after missing their first, although this delays benefit receipt). The downstream outcome $Y\in\mathbb{R}^{+}$ that we consider is the cumulative total of benefits received over the next 12 months. Missing an interview delays benefit receipt, but Homonoff and Somerville (2021) find that outcomes are skewed: Some would-be recipients do not further pursue recertification after missing an interview, potentially losing out on thousands of dollars’ worth of benefits. Homonoff and Somerville (2021) establish via linkage to administrative wage data that many of these individuals are indeed SNAP-eligible and thus they rule out explanations based on purely economic models of ordeal targeting, which interpret a beneficiary’s missing an interview to mean that her household does not need benefits. This finding indicates that behavioral frictions in service delivery can have negative effects. In other words, reduction of frictions can improve service delivery.
Next, we include descriptive statistics to provide evidence on the existence of treatment effect and compliance heterogeneity and potential drivers of inequity in unconstrained optimal encouragement regimes by disaggregating the treatment effect and compliance heterogeneity (vs. the heterogeneity in the treatment effect of the nudge alone). In LABEL:tbl-regressionRX, we display the results of a logistic regression of recommendation on covariates. In the original data, significant explanatory factors include household size, whether the interview was conducted by phone or in person, gender (female), race (nonwhite), age, whether the beneficiary has children, whether the beneficiary receives the maximum benefit amount, and time lapse to the interview date. Next, in LABEL:tbl-regressionTRX, we include the regression table from a recommendation-interacted logistic regression of treatment $T$ on $(X,R,R\times X),$ i.e., covariates and interacted interview reminder. We find significant heterogeneity based on age (interview reminders are less efficacious as age increases), while reminder efficacy increases for those without income in the previous quarter (which indicates more need for benefits/a lower opportunity cost of interview attendance) and with years since first SNAP enrollment (which indicates greater familiarity with the recertification process).
In Figure 2, we compare the distributions within each race (binarized to white or nonwhite) of the heterogeneous treatment effect $\mathbb{E}[Y(1)-Y(0)\mid X],$ the heterogeneous compliance effect $p_{1\mid 1}(X)-p_{1\mid 0}(X),$ and the heterogeneous encouragement effect, which is the product of the previous two: $(\mathbb{E}[Y(1)-Y(0)\mid X])(p_{1\mid 1}(X)-p_{1\mid 0}(X))$. In a setting with fixed costs of recommendation and treatment, the default approach of viewing the recommendation as the treatment, i.e., an optimal unconstrained encouragement regime, sorts by the latter score.
In this setting, we find that treatment efficacy is broadly correlated with takeup: Higher treatment effects are rank correlated with higher compliance effects. However, the strength of this relationship varies across groups, leading to inequality. We quantify this via the Spearman’s $\rho$ rank correlation coefficient. It measures the correlation of the ranks of two random variables and hence indicates whether a monotonic (but not necessarily linear) relationship exists between them. Among white beneficiaries only, the rank correlation coefficient between heterogeneous treatment and compliance effects is 0.68, while it is 0.45 for nonwhite beneficiaries. While both correlations are statistically significantly positive, the relationship is weaker for nonwhite than for white beneficiaries, in particular with regard to those most likely to take up. This differential efficacy can lead to potential inequity.
A central concern here is that worse-off groups might have more to gain by accessing the treatment (i.e., have heterogeneous treatment effects of greater magnitude) but may respond less to nudges/recommendations into treatment (because of behavioral frictions, time/resource constraints, higher need, etc.). In this setting, decreasing marginal utilities of money establish the former, while the process of filling out the burdensome benefits application or making it to a recertification interview under time, childcare, transportation, or resource constraints speaks to the relative difficulty posed by the latter. The density plots in Figure 2 underscore this phenomenon. The first plot shows the treatment effect heterogeneity: The distribution is wider for nonwhite than for white beneficiaries; i.e., it has more mass at higher magnitudes of $\tau$ and less at lower magnitudes of $\tau$. However, we see the opposite effect on compliance: While the results are relatively weak, they are bimodal, with a substantially greater probability density at highercompliance treatment effects for white than for nonwhite beneficiaries. One reason for this could be the language barrier for non-English-speaking applicants.
The last plot shows the heterogeneity in the encouragement effect: When we combine the treatment and compliance effects, we see that the disparities in the compliance effects result in the distribution of the heterogeneous encouragement effect for white beneficiaries being wider than that for nonwhite ones. On the basis of such results, resource-constrained optimal rules would tend to target outreach to white individuals because of their higher “nudgeability". This could be concerning because, in some sense, the efficacy of recommendations can be changed by means of tactical redesign of the recommendation: more intensive outreach, pairing with a benefits counselor, etc. Koenecke et al. (2023) study efficient vs. equitable budget allocation for SNAP outreach via public surveys and find broad public support for equitable allocations, with potentially higher spending for targeting of harder-to-reach minority groups, for example.
7.1.3 Comparison of potential improvements upon the status quo
$p_{1\mid 1}(x,a)-p_{1\mid 0}(x,a)$ $\tau(x)$ $P(T=1\mid X,A)$ $E[Y\mid X,A]$
Correlation -0.08 -0.28 -0.27 0.04
Correlation A=0 0.23 -0.27 -0.21 0.21
Correlation A=1 -0.15 -0.28 -0.27 0.04
Table 1: Self-selection and targeting efficiency: Spearman’s rank correlation of predicted enrollment in reminder probabilities ($P(R=1\mid X,A)$) with heterogeneous and marginal compliance and treatment effects.
Is self-selection sufficient for targeting?
An important question is whether individuals self-select into the reminder on the basis of their expected value of benefits or the particular effect of the reminder on compliance. We leverage the estimated effect heterogeneity and predictive models to unpack this relationship. We consider the Spearman correlation between the predicted probability of enrollment in the reminder, $P(R=1\mid X,A)$, and $p_{1\mid 1}(x,a)-p_{1\mid 0}(x,a)$ (the heterogeneous effect of compliance); $\tau(X)$ (the heterogeneous effect of treatment), $P(T=1\mid X,A)$ (the marginal treatment-enrollment probabilities); and $E[Y\mid X,A]$ (the marginal expected benefits). The first two reflect the effect heterogeneity: A strong positive rank correlation of selection into encouragement (the reminder) with the compliance or heterogeneous treatment effects (or the product thereof) would indicate that self-selection is efficient for recommendation/treatment efficacy. Since this may be difficult to justify a priori, as enrolling individuals do not know this estimated heterogeneity in the effects, we also assess rank correlations of $P(R=1\mid X,A)$ with the marginal treatment and outcomes. We find that, overall, self-selection is weakly negatively correlated with heterogeneous compliance, treatment effects, and marginal compliance. However, these relationships are heterogeneous by group membership: We find a weakly positive correlation for the white subgroup, where self-selection is weakly correlated with heterogeneous compliance effects and marginal outcome levels. However, generally no such relationship exists for nonwhite beneficiaries. This indicates that current self-selection patterns are overall not efficient for the targeting of reminders to impact outcomes and that, in fact, self-selection is usually negatively correlated with the heterogeneity in the treatment effect: Better-resourced households with lower treatment effects (eligible for lower benefits amounts) are somewhat more likely to enroll in recertification reminders. At a high level, this suggests that there is room for improvement upon self-selection into deadline reminders, whether by changing the default option to a reminder opt-out or by means of covariate-conditional targeting under a budget constraint, as we explore next. A recent Federal Communications Commission (FCC) ruling (Federal Communications Commission, 2021) clarifies that “state governments can communicate information to beneficiaries via text messaging without first obtaining their ‘opt-in’ and without facing liability under the TCPA [Telephone Consumer Protection Act of 1991]”.
Comparison of automatic enrollment in reminders (recommend to all) and optimized budgeted allocation
Next, we consider potential improvements from automatic enrollment in reminders (recommend to all) or, under potential budget constraints, budget-optimal targeted allocations. In Figure 1, we compare all of these potential policies. The y-axis measures different values of $E[Y(\pi)]$ in differently colored series of lines (red, purple, blue): red for the group $A=1$ (nonwhite), $E[Y(\pi)\mid A=1]$; purple for the entire population, $E[Y(\pi)]$; and blue for $A=0$ (white), $E[Y(\pi)\mid A=0]$. We optimize the policies on a 70% training split. The off-policy values are obtained via doubly robust estimation, with logistic regression as the propensity score and gradient-boosted regression on log-transformed outcomes for outcome regression. However, because of the zero-inflated/heavy-tailed outcomes, we find some instability in the off-policy estimates computed on the test set alone. Thus, we pool the data and obtain the plotted estimates of $E[Y(\pi)\mid A=1]$ from the entire dataset.
Next, within each set of colored lines, we compare the following: The solid line ranging over budgets indicates $\mathbb{E}[Y(\pi_{\%})]$, where $\pi_{\%}$ is the optimal policy under a budget corresponding to delivery of recommendations to some percentage of the population. The x-axis shows this percentage ranging from 0% (no one is treated) to 100% (everyone is treated). Next, the set of dotted/dashed lines presents the outcomes under unpersonalized or untargeted policies: The densely dotted line indicates treat-all, $E[Y(1)\mid A=a]$; the dashed line indicates the observed self-selection (and hence averages over outcomes in the dataset within group status); and the sparsely dotted line indicates $E[Y(1)\mid A=a]$. Within a set of colored lines, on the left side, the budget is large, so the value of the targeted allocation, $E[Y(\pi_{\%})]$ (solid line), tracks the value of therecommend-to-all policy, $E[Y(1)]$; while on the right side, the value of small budgets tracks the value of the recommend-to-none policy, $E[Y(0)]$.
First, we note that, for the most part, any additional reminders generally improve outcomes—whether via self-selection, targeted allocation, or recommendation-to-all. However, as we saw in the previous analysis of self-selection, the value under currently observed self-selection does not improve significantly because of inefficient targeting, although there is heterogeneity: Self-selection somewhat improves outcomes for the white group. Of course, this depends on many factors, some of which may shift in different settings and implementations. Race may be correlated with income and digital literacy (enrollment with phone number), and the reminders differed by device. For most budgets covering treatment of at least $25\%$ of the population, we find that targeted allocation improves group-average outcomes significantly over those under current self-selection or all-control (no reminders). To conclude, regarding equity in outcomes, i.e., group-level outcomes $E[Y(1)\mid A=1],E[Y(1)\mid A=0]$, targeted allocation improves upon a self-selection or no-recommendation status quo.
Figure 1: Comparison of average group outcomes under budget allocations for targeted treatment of different beneficiary shares ($E[Y(\pi_{\%})\mid A=a]$) with self-selection ($E[Y(\pi_{0})\mid A=a]$) or no-reminder ($E[Y(0)\mid A=a]$) status quo.
Figure 2: SNAP recertification case study: Heterogeneous treatment effects $\tau$, compliance, and their product (heterogeneous encouragement effects), density distribution plots by race.
Figure 3: SNAP recertification case study. Each figure indicates the performance metric (conditional encouragement effect, i.e., compliance score $\times$ heterogeneous treatment effect; heterogeneous treatment effect; and compliance score induced by a threshold policy that thresholds on $\left(p_{1\mid 1}-p_{1\mid 0}\right)\times\tau$).
Our final analysis considers equity in allocation efficiency, specifically the lift or improvements to decisions around who receives recommendations. In Figure 3, we evaluate the lift in different performance metrics conditional on protected group and conditional on the group’s being encouraged under a resource-constrained policy—the y-axes show $\mathbb{E}[c(Y\pi))\mid A=a,\pi(X)=1],\mathbb{E}[(Y(1)-Y(0))\mid A=a,\pi(X)=1],\mathbb{E}[(T(1)-T(0))\mid A=a,\pi(X)=1]$. The x-axis ranges over thresholds on the conditional encouragement effect, $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\times\tau(X,A)$. Thus, the x-axis explores the range of the population covered under the budget allocation, with the budget averaged over the full population on the left-hand side and encouragement restricted to few individuals on the right-hand side. This represents the range of resource-constrained rules under a naive ITT approach. We do find, for large ranges of moderate budgets, that the naive efficient allocation results in disparities in average outcomes for white and nonwhite beneficiaries. The next two figures on the right disaggregate the individual encouragement effect into the heterogeneous causal effect $\tau$ of interview attendance and the compliance effect $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))$ of text message reminders on interview attendance. Consistent with the earlier results on the strong positive rank correlation of $\tau$ and $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))$, both plots are nearly nondecreasing as we threshold based on increasing individual encouragement scores $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\times\tau$. However, as we also saw earlier, this correlation is weaker for nonwhite beneficiaries, and we can observe a small region of nonmonotonic behavior. Comparing these two plots, we see that the disparities in naive ITT-based allocation are in fact driven almost entirely by disparities in access, which would persist under a large range of budgets. Disaggregating our analysis by looking separately at treatment vs. compliance effects and at covariate-heterogeneous effects allows us to quantify access gaps that would persist even under efficient allocations.
Figure 4: Groupwise separate budgets: Which budget allocations would achieve equal improvements?
Table 2: Equivalent budgets for $A=1$ to equalize improvement at a budget for $A=0$
Budget A=0 Improvement A=0 Equiv. Budget A=1
50.0% \$12.22 71.7%
25.0% \$21.97 50.2%
10.0% \$39.37 29.2%
With these findings in mind, we turn to a final investigation: Suppose that we could consider allocations with different budgets per group rather than a global budget. Given that different compliance responses lead to unequal improvements, how much more outreach toward nonwhite populations would be needed under the allocation to equalize the amount of improvement over the no-allocation baseline? We analyze this question in Figure 4. We now examine policies that consider different thresholds for different groups, thereby resulting in different budgets, $\pi_{\%,A}=\mathbb{I}[(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)>q_{A}].$ Typically, $q_{A}$ is a quantile of the score distribution and therefore corresponds to some budget percentile. We let $F_{A}(p)=P(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)\leq p)$ denote the cumulative distribution function (CDF) of the heterogeneous encouragement score, and therefore, the groupwise budget at some threshold $q_{a}$ is $F^{-1}_{A}(q_{A})$. We now range over these thresholds separately for the different groups on the $x$-axis; the $y$-axis plots the average improvement in the allocated subgroup, $E[p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)\mid A,\pi_{\%,A}=1]$. In Table 2, we illustrate for a few potential groupwise budget percentages of $A=0$ what improvement is achieved for $A=0$ and what percentage of the budget would be required for $A=1$ to achieve equal improvement. The equivalent budget allocations required to target the same percentage of the group population can differ by 20 to 25 percentage points. At thresholds $q=F_{0}^{-1}(p)$ corresponding to percentiles $p=50\%,75\%$ and $90\%$ of the $A=0$ group, and therefore budgets of $1-p=50\%,25\%$ and $10\%$, we observe average improvements of \$$\$12,\$21.97$ and \$$\$39.37$ in the $A=0$ group, and it would require an equivalent budget of ${}72\%,50\%$ and $29.2\%$ of the $A=1$ population to achieve equal improvements, respectively. Given that 80% of individuals are nonwhite (i.e., the $A=1$ group is four times the size of the $A=0$ group), substantially different groupwise budgets (in terms of number of people covered by the budget allocation) would be required to equalize improvements.
Summary and Conclusion
In this case study, we focus on assessing the heterogeneous compliance and treatment effects of text message reminders about recertification deadlines on 1) completion of any interview and 2) yearly benefits. First, we focus on insights on the self-selection process and heterogeneity in the encouragement and treatment effects that help us understand potential drivers of inequity in targeted allocations. Although we find that the heterogeneity in the effects of the encouragement (text message reminders about the recertification deadline) on treatment (successful completion of an interview) is positively correlated with the heterogeneity in the treatment effects on downstream benefits collected, the strength of this correlation varies by race: It is stronger for white than for nonwhite beneficiaries. This finding is relevant since the naive heterogeneous encouragement effect can be related to the product of heterogeneous encouragement and treatment effects and so the relative strength of this correlation indicates that an approach based purely on ranking the heterogeneous encouragement effects under the efficient budget allocation would result in some white beneficiaries ranking more highly than nonwhite ones.
Figure 2 and Figure 3 illustrate that this inequity arises because of the differential effectiveness of the reminder in improving the probability of interview attendance, suggesting either that a design with further communication could be helpful or that other, structural compliance barriers could persist. Proceeding with the analysis of reminder efficacy, we find that while racial disparities in current patterns of self-selection into encouragement are small, self-selection is still inefficient for targeting: Those who would benefit the most from receiving reminders do not sign up for them. We assess budget-optimal targeted allocations that can improve upon the outcome under self-selection with as little as 10% of the budget. The heterogeneous effects of the reminder on interview completion, i.e., the differences that the nudge makes, are smaller for nonwhite beneficiaries, but this group generally receives higher benefits overall. Overall, any efficiency-targeted allocation rule improves groupwise outcomes (i.e., average 12-month benefits for white or nonwhite beneficiaries) relative to those in the no-reminder baseline.
Our overall findings highlight the need for further research into the roots of the gaps in interview completion and suggest that investing in encouragement could reduce racial gaps and may be overall more effective than constraining allocations for fairness. We consider a hypothetical fairness adjustment in which we vary the encouragement budgets for groups or the thresholds at which groups receive the encouragement, finding that substantially different budgets across groups would be required to equalize improvements—although this could also be perceived as unfair. Overall, we find that naively encouragement-optimal allocations balance fair treatment (less inequality in budget spending) with equity in outcomes.
In conclusion, we find evidence of effect heterogeneity that implies that targeted allocations or opt-out reminders can improve significantly upon the outcomes under current self-selection. Text message reminders and communication infrastructure open up a tactical space for operational service design. Investigating potential targeted allocations, we find that they improve equity in final outcome levels but that inequity persists in how much groups’ outcomes improve on average, driven almost entirely by racial gaps in compliance. Although our tools enable us to dig deeper into the efficacy of self-selection and potential alternative allocations and estimate disparities in outcomes, more social science research is necessary to elucidate the root causes. [^1] Commentaries on nudges in other domains (e.g., criminal justice Kohler-Hausmann, 2020) highlight varying explanations consistent with nonresponse. This is especially important because whether the observed disparities in compliance reflect behavioral inattention, trust in government, or structural barriers to compliance with administrative burdens matters for policy implications.
7.2 Case study: Decision-making framework for electronic monitoring.
We conduct a case study on a dataset of judicial decisions on supervised release based on risk-score-informed recommendations for supervised release under an electronic-monitoring program (Office of the Chief Judge, 2019a). The Public Safety Assessment Decision-Making Framework (PSA-DMF) uses a prediction of failure to appear for a future court date to inform pretrial decisions, including the decision that we focus on here: supervised release with electronic monitoring, where judges make the final decision (psa, 2016). While the literature on pretrial risk assessment is large, to our knowledge, it is unclear what empirical evidence justifies the release recommendation matrices that have been used in practice to recommend supervised release via electronic monitoring. [^2] We focus on supervised release with electronic monitoring, though the broad term supervised release encompasses substantially different programs nationwide, including access to supportive services and caseworkers, which has been touted as a factor enabling bail reform and release more broadly (Akinnibi and Holder, 2023). There are current policy concerns about disparities in the increasing use of supervised release, given mixed evidence on its outcomes (Office of the Chief Judge, 2019a; Gross, ); e.g., Safety and Justice Challenge (2022) concludes that "targeted efforts to reduce racial disparities [in supervising release] are necessary". We focus on a publicly available dataset from Cook County, Illinois, with information about defendant characteristics, algorithmic recommendations for electronic monitoring, detention/release/supervised release decisions, and failure to appear and other outcomes (Office of the Chief Judge, 2019b). The data were initially used to assess bail reform (Office of the Chief Judge, 2019a). [^3] While in this initial case study we work with publicly available data (Office of the Chief Judge, 2019b), in future work, we will seek more granular data with additional robustness checks to support our substantive conclusions. We offer a more detailed discussion in the appendix, but to summarize, unconfoundedness is likely violated in this case (but can be addressed with standard sensitivity analysis), and some line-level data are aggregated for privacy.
We let $Z\in\{0,1\}$ denote release $(Z=1)$ (with or without conditions). All of our analysis occurs in the $(XZ,AZ,RZ,YZ)$ group, i.e., among the released population only. For brevity, we drop $Z$ in describing the data below. We let $X$ denote covariates (age, top charge category, PSA failure to appear/new criminal arrest (FTA/NCA) score bucket and flag). The (binarized) protected attribute is $A$: race (nonwhite/white) or gender (female/male). The algorithmic recommendation is $R$, a recommendation from the PSA-DMF matrix for supervised release (at any intensity of supervision conditions). The treatment $T$ is whether the individual is released under supervision (at any intensity of supervision conditions). The outcome variable, $Y$, is failure to appear ($Y=1$).
In Figure 6, we provide descriptive information illustrating heterogeneity (including by protected attribute) in adherence and effectiveness. We observe wide variation in judges’ assignment of supervised release beyond the recommendation. We use logistic regression to estimate outcome and treatment response models. The first figure shows our estimates of the causal effect by gender (with similar heterogeneity by race). The outcome is failure to appear, so negative scores are beneficial. The second figure illustrates the difference in responsiveness: how much more likely decision-makers are to assign treatment when there is than when there is not an algorithmic recommendation to do so. The last figure plots a logistic regression of the lift in responsiveness on the causal effect $\tau(x,a)=\mu_{1}(x,a)-\mu_{0}(x,a)$. We observe disparities in how responsive decision-makers are conditional on the same treatment effect efficacy. This is, importantly, not a claim of animus because decision-makers did not have access to causal effect estimates. Nonetheless, disparities persist.
Figure 5: Distribution of treatment effect by gender, lift in treatment probabilities $p_{11a}-p_{01a}=P(T=1\mid R=1,A=a,X)-P(T=1\mid R=0,A=a,X)$, and plot of $p_{11a}-p_{01a}$ vs. $\tau.$
Figure 6: Policy value $V(\pi^{\lambda})$, treatment value $\mathbb{E}[T(\pi^{\lambda})\mid A=a],$ for $A$ = race, gender.
In Figure 6, we highlight results from constrained policy optimization. The first two plots in each set illustrate the objective function value and $A=a$ average treatment cost for $A=$ race (nonwhite/white) and gender (female/male), respectively. We use costs of $100$ for $Y=1$ (failure to appear, 00 for $Y=0$, and $20$ when $T=1$ (set arbitrarily)) and minimize costs. On the x-axis, we plot the penalty $\lambda$ that we use to assess the solutions of Proposition 9. The vertical dashed line indicates the solution achieving $\epsilon=0$, i.e., parity in treatment takeup. Near-optimal policies that reduce treatment disparity can be of interest given advocacy concerns about how the expansion of supervised release could increase the surveillance of already surveillance-burdened marginalized populations. We see that, indeed, for race, surveillance parity–constrained policies can substantially reduce disparities for nonwhite defendants while not increasing surveillance on white defendants that much: The red line decreases significantly with a low increase of the blue line (and low increases to the objective value). On the other hand, for gender, the opportunity for improvement in the surveillance disparity is much smaller.
Acknowledgments
We thank the San Francisco Human Services Agency for access to the data and Peri Weisberg for data access and discussions.
Appendix A Additional discussion
A.1 Additional discussion on motivating fairness constraints
[Christensen et al., 2020] posits a “human capital catch-22" that certain axes of precarity such as scarcity and health both increase likelihood of requiring access to state assistance while reducing the cognitive resources required to navigate administrative burdens in service delivery. Some strategies noted for reducing administrative burden in public benefit and service programs OMB [2022] include reducing information/learning costs, which can be modeled with encouragement designs and targeted recommendations; or reducing redemption costs, which can be modeled in a setting of scarce resources with constraints and potentially personalized using the methods we develop here.
A.2 Additional related work
Intention-to-treat analysis. We appeal to intention-to-treat analysis with randomness that either arises from human decision-makers or individual non-adherence/non-compliance, but we generally assume the data does not include information about the identity of different decision-makers, which is common with publicly available data. Our conditional exclusion restriction also means that certain decomposed effects are zero, so mediation analysis is less relevant. A related literature studies principal stratification Jiang et al. [2020], which is less interpretable since stratum membership is unknown. Similarly, even though encouragement effects are driven by compliers, complier-conditional analysis is less policy-relevant since complier identities are unknown. In general, our causal identification arguments are based on covariate-adjusted intention-to-treat analysis and covariate-adjusted as-treated analysis. We avoid estimation of stratum-specific effects, because if complier status is unknown, prescriptive decision rules cannot directly personalize by stratum membership.
Fair off-policy learning. We highlight some most closely related works in off-policy learning (omitting works in the sequential setting). [Metevier et al., 2019] studies high-probability fairness constraint satisfaction. [Kim et al., 2022] studies doubly-robust causal fair classification, while others have imposed deterministic resource constraints on the optimal policy formulation [Chohlas-Wood et al., 2021]. Other works study causal or counterfactual risk assessments [Mishler et al., 2021, Coston et al., 2020]. Our perspective is closer to that of off-policy learning, i.e. approximating direct control over the intervention by assuming stability in decision-maker treatment assignment probabilities. [Kallus and Zhou, 2019] studies (robust) bounds for treatment responders in binary outcome settings; this desiderata is coupled to classification notions of direct treatment. Again, our focus is on modeling the fairness implications of non-adherence. Indeed, in order to provide general algorithms and methods, we do build on prior fair classification literature. A different line of work studies “counterfactual" risk assessments which models a different concern.
Principal stratification and mediation analysis in causal inference
[Liu et al., 2021] studies an optimal test-and-treat regime under a no-direct-effect assumption, that assigning a diagnostic test has no effect on outcomes except via propensity to treat, and studies semiparametric efficiency using Structural Nested-Mean Models. Though our exclusion restriction is also a no-direct-effect assumption, our optimal treatment regime is in the space of recommendations only as we do not have control over the final decision-maker, and we consider generally nonparametric models.
We briefly go into more detail about formal differences, due to our specific assumptions, that delineate the differences to mediation analysis. Namely, our conditional exclusion restriction implies that $Y_{1T_{0}}=Y_{T_{0}}$ and that $Y_{0T_{1}}=Y_{1T_{1}}$ (in mediation notation with $T_{r}=T(r)$ in our notation), so that so-called net direct effects are identically zero and the net indirect effect is the treatment effect (also called average encouragement effect here).
Human-in-the-loop in consequential domains. There is a great deal of interest in designing algorithms for the “human in the loop" and studying expertise and discretion in human oversight in consequential domains [De-Arteaga et al., 2020]. On the algorithmic side, recent work focuses on frameworks for learning to defer or human-algorithm collaboration. Our focus is prior to the design of these procedures for improved human-algorithm collaboration: we primarily hold fixed current human responsiveness to algorithmic recommendations. Therefore, our method can be helpful for optimizing local nudges. Incorporating these algorithmic design ideas would be interesting directions for future work.
Empirical literature on judicial discretion in the pretrial setting.
Studying a slightly different substantive question, namely causal effects of pretrial decisions on later outcomes, a line of work uses individual judge decision-makers as a leniency instrumental variable for the treatment effect of (for example, EM) on pretrial outcomes [Arnold et al., 2022, 2018, Lum et al., 2017]. And, judge IVs rely on quasi-random assignment of individual judges. We focus on the prescriptive question of optimal recommendation rules in view of patterns of judicial discretion, rather than the descriptive question of causal impacts of detention on downstream outcomes.
A number of works have emphasized the role of judicial discretion in pretrial risk assessments in particular [Green and Chen, 2021, Doleac and Stevenson, 2020, Ludwig and Mullainathan, 2021]. In contrast to these works, we focus on studying decisions about electronic monitoring, which is an intermediate degree of decision lever to prevent FTA that nonetheless imposes costs. [Imai et al., 2020] study a randomized experiment of provision of the PSA and estimate (the sign of) principal causal effects, including potential group-conditional disparities. They are interested in a causal effect on the principal stratum of those marginal defendants who would not commit a new crime if recommended for detention. [Ben-Michael et al., 2021] study policy learning in the absence of positivity (since the PSA is a deterministic function of covariates) and consider a case study on determining optimal recommendation/detention decisions; however their observed outcomes are downstream of judicial decision-making. Relative to their approach, we handle lack of overlap via an exclusion restriction so that we only require ambiguity on treatment responsivity models rather than causal outcome models.
Appendix B Additional discussion on method
B.1 Extrapolation in 2
A naive approach based on parametric extrapolation estimates $p_{1\mid 1}(X)$, treatment responsivity, on the observed data and simply uses the parametric form to extrapolate to the full dataset. In the case study later on, the support of $X\mid R=1$ is a superset of the support of $X\mid R=0$ in the observational data. Given this, we derive the following alternative identification based on marginal control variates (where $p_{t}=P(T=t\mid X)$ marginalizes over the distribution of $R$ in the observational data):
Proposition 6 (Control variate for alternative identification ).
Assume $Y(T(r))\perp T(r)\mid R=r,X$.
$\textstyle V(\pi)=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\left\{c_{rt}(Y(t))\frac{\mathbb{I}\left[T=t\right]}{p_{t}(X)}+\left(1-\frac{\mathbb{I}\left[T=t\right]}{p_{t}(X)}\right)\mu_{t}(X)\right\}p_{t\mid r}(X)\right]$
On the other hand, parametric extrapolation is generally unsatisfactory because conclusions will be driven by model specification rather than observed data. Nonetheless, it can provide a starting point for robust extrapolation of structurally plausible treatment response probabilities.
Appendix C Additional discussion on general optimization method
C.1 Additional discussion on constrained optimization
Feasibility program
We can obtain upper/lower bounds on $\epsilon$ in order to obtain a feasible region for $\epsilon$ by solving the below optimization over maximal/minimal values of the constraint:
$\overline{\epsilon},\underline{\epsilon}\in\max_{\pi}/\min_{\pi}\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]$ (2)
$V_{\epsilon}^{*}=\max_{\pi}\;\{\mathbb{E}[c(\pi,T(\pi),Y(\pi))]\colon\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]\leq\epsilon\}$ (3)
C.2 Additional discussion on Algorithm 2 (general algorithm)
C.2.1 Weighted classification reduction and off-policy estimation
There is a well-known reduction of optimizing the zero-one loss for policy learning to weighted classification. Note that the reductions approach of [Agarwal et al., 2014] works with the Lagrangian relaxation which only further introduces datapoint-dependent additional weights. Notationally, in this section, for policy optimization, $\pi\in\{-1,+1\},T\in\{-1,+1\}$ (for notational convenience alone). We consider parameterized policy classes so that $\pi(x)=\pi(1\mid x)=\operatorname{sign}(f_{\beta}(x))$ for some index function $f$ depending on a parameter $\beta\in\mathbb{R}^{d}$. Consider the centered regret $J(\pi)=\mathbb{E}[Y(\pi)]-\frac{1}{2}\mathbb{E}[\mathbb{E}[Y\mid R=1,X]+\mathbb{E}[Y\mid R=0,X]].$ We summarize different estimation strategies via the score function $\psi_{(\cdot)}(O),$ where ()$(\cdot)\in\{DM,IPW,DR\}$: the necessary property is that ${\mathbb{E}[\psi\mid X]=\mathbb{E}[Y\mid R=1,X]-\mathbb{E}[Y\mid R=0,X]}$. The specific functional forms of these different estimators are as follows, where $\mu^{R}_{r}(X)=\mathbb{E}[Y\mid R=r,X]:$
$\psi_{DM}=(p_{1\mid 1}(X)-p_{1\mid 0}(X))(\mu_{1}(X)-\mu_{0}(X)),\psi_{IPW}=\frac{RY}{e_{R}(X)},\psi_{DR}=\psi_{DM}+\psi_{IPW}+\frac{R\mu^{R}(X)}{e_{R}(X)}.$
C.2.2 Additional fairness constraints and examples in this framework
In this section we discuss additional fairness constraints and how to formulate them in the generic framework. Much of this discussion is quite similar to [Agarwal et al., 2018] (including in notation) and is included in this appendix for completeness only. We only additionally provide novel identification results for another fairness measure on causal policies in Section C.2.3, concrete discussion of the reduction to weighted classification, and provide concrete descriptions of the causal fairness constraints in the more general framework.
We first discuss how to impose the treatment parity constraint. This is similar to the demographic parity example in Agarwal et al. [2018], with different coefficients, but included for completeness. (Instead, recommendation parity in $\mathbb{E}[\pi\mid A=a]$ is indeed nearly identical to demographic parity.)
Example 5 (Writing treatment parity in the general constrained classification framework.).
We write the constraint
$\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]$ (4)
in this framework as follows:
$\mathbb{E}[T(\pi)\mid A=a]=\mathbb{E}[\pi_{1}(X)(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))+p_{1\mid 0}(X,A)\mid A=a]$
For each $u\in\mathcal{A}$ we enforce that
$\sum_{r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)p_{1\mid r}(X,A)\mid A=u\right]=\sum_{r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X,A)p_{1\mid r}(X,A)\right]$
We can write this in the generic notation given previously by letting $\mathcal{J}={\mathcal{A}\cup\{\circ\}},$
$g_{j}(O,\pi(X);\eta)=\pi_{1}(X)(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))+p_{1\mid 0}(X,A),\forall j.$
We let the conditioning events $\mathcal{E}_{a}=\{A=a\},\mathcal{E}_{\circ}=\{\text{True}\},$ i.e. conditioning on the latter is equivalent to evaluating the marginal expectation. Then we express Equation 4 as a set of equality constraints $h_{a}(\pi)=h_{\circ}(\pi),$ leading to pairs of inequality constraints,
$\left\{\begin{matrix}h_{u}(\pi)-h_{\circ}(\pi)\leq 0\\ h_{\circ}(\pi)-h_{u}(\pi)\leq 0\end{matrix}\right\}_{u\in\mathcal{A}}$
[…]
Next, we discuss a more complicated fairness measure. We first discuss identification and estimation before we also describe how to incorporate it in the generic framework.
C.2.3 Responder-dependent fairness measures
We consider a responder framework on outcomes (under our conditional exclusion restriction). Because the contribution to the AEE is indeed from the responder strata, this corresponds to additional estimation of the responder stratum.
We enumerate the four possible realizations of potential outcomes (given any fixed recommendation) as $(Y(0(r)),Y(1(r))\in$$\{0,1\}^{2}$. We call units with $(Y(0(r)),Y(1(r)))=(0,1)$ responders, $(Y(0(r)),Y(1(r)))=(1,0)$ anti-responders, and $Y(0(r))=Y(1(r))$ non-responders. Such a decomposition is general for the binary setting.
Assumption 9 (Binary outcomes, treatment).
$T,Y\in\{0,1\}$
Assumption 10 (Monotonicity).
$Y(T(1))\geq Y(T(0))$
Importantly, the conditional exclusion restriction of Assumption 2 implies that responder status is independent of recommendation. Conditional on observables, whether a particular individual is a responder is independent of whether someone decides to treat them when recommended. In this way, we study responder status analogous to its use elsewhere in disparity assessment in algorithmic fairness [Imai et al., 2020, Kallus et al., 2019a]. Importantly, this assumption implies that the conditioning event (of being a responder) is therefore independent of the policy $\pi$, so it can be handled in the same framework. s
We may consider reducing disparities in resource expenditure given responder status.
We may be interested in the probability of receiving treatment assignment given responder status.
Example 6 (Fair treatment expenditure given responder status).
$\mathbb{E}[T(\pi)\mid Y(1(R))>Y(0(R)),A=a]-\mathbb{E}[T(\pi)\mid Y(1(R))>Y(0(R)),A=b]\leq\epsilon$
We can obtain identification via regression adjustment:
Proposition 7 (Identification of treatment expenditure given responder status).
Assume Assumptions 9 and 10.
$P(T(\pi)=1\mid A=a,Y(1(\pi))>Y(0(\pi)))=\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]}$
Therefore this can be expressed in the general framework.
Example 7 (Writing treatment responder-conditional parity in the general constrained classification framework.).
For each $u\in\mathcal{A}$ we enforce that
$\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=u])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=u]}=\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))]}$
We can write this in the generic notation given previously by letting $\mathcal{J}={\mathcal{A}\cup\{\circ\}},$
$g_{j}(O,\pi(X);\eta)=\frac{\{\pi_{1}(X)(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))+p_{1\mid 0}(X,A)\}(\mu_{1}(X,A)-\mu_{0}(X,A))}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]},\forall j.$
Let $\mathcal{E}_{a}^{j}=\{A=a_{j}\},\mathcal{E}_{\circ}=\{\text{True}\},$ and we express Equation 4 as a set of equality constraints of the above moment $h_{a}(\pi)=h_{\circ}(\pi),$ leading to pairs of inequality constraints,
$\left\{\begin{matrix}h_{u}(\pi)-h_{\circ}(\pi)\leq 0\\ h_{\circ}(\pi)-h_{u}(\pi)\leq 0\end{matrix}\right\}_{u\in\mathcal{A}}$
The corresponding coefficients of $M$ proceed analogously as for treatment parity.
C.2.4 Best-response oracles
Best-responding classifier $\pi$, given $\lambda$: $\operatorname{BEST}_{\pi}(\lambda)$
The best-response oracle, given a particular $\lambda$ value, optimizes the Lagrangian given $\pi$:
[…]
Best-responding Lagrange multiplier $\lambda$, given $\pi$:
$\operatorname{BEST}_{\boldsymbol{\lambda}}(Q)$ is the best response of the $\Lambda$ player. It can be chosen to be either 00 or put all the mass on the most violated constraint. Let ${{\gamma}}(Q):={M}{{h}}(Q)$ denote the constraint values, then $\operatorname{BEST}_{\boldsymbol{\lambda}}(Q)$ returns
$\begin{cases}\mathbf{0}&\text{ if }\widehat{\gamma}(Q)\leq\widehat{\mathbf{c}}\\ B\mathbf{e}_{k^{*}}&\text{ otherwise, where }k^{*}=\arg\max_{k}\left[\widehat{\gamma}_{k}(Q)-\widehat{c}_{k}\right]\end{cases}$
C.2.5 Weighted classification reduction
There is a well-known reduction of optimizing the zero-one loss for policy learning to weighted classification. A cost-sensitive classification problem is
$\underset{\pi_{1}}{\arg\min}\sum_{i=1}^{n}\pi_{1}\left(X_{i}\right)C_{i}^{1}+\left(1-\pi_{1}\left(X_{i}\right)\right)C_{i}^{0}$
The weighted classification error is $\sum_{i=1}^{n}W_{i}{1}\left\{h\left(X_{i}\right)\neq Y_{i}\right\}$ which is an equivalent formulation if $W_{i}=\left|C_{i}^{0}-C_{i}^{1}\right|$ and $Y_{i}={1}\left\{C_{i}^{0}\geq C_{i}^{1}\right\}$.
The reduction to weighted classification is particularly helpful since taking the Lagrangian will introduce datapoint-dependent penalties that can be interpreted as additional weights. We can consider the centered regret $J(\pi)=\mathbb{E}[Y(\pi)]-\frac{1}{2}\mathbb{E}[\mathbb{E}[Y\mid R=1,X]+\mathbb{E}[Y\mid R=0,X]]$. Then
$J(\theta)=J(\operatorname{sgn}(g_{\theta}(\cdot)))=\mathbb{E}[\operatorname{sgn}(g_{\theta}(X))\left\{\psi\right\}]$
where $\psi$ can be one of, where $\mu^{R}_{r}(X)=\mathbb{E}[Y\mid R=r,X],$
$\psi_{DM}=(p_{1\mid 1}(X)-p_{1\mid 0}(X))(\mu_{1}(X)-\mu_{0}(X)),\psi_{IPW}=\frac{RY}{e_{R}(X)},\psi_{DR}=\psi_{DM}+\psi_{IPW}+\frac{R\mu^{R}(X)}{e_{R}(X)}$
We can apply the standard reduction to cost-sensitive classification since $\psi_{i}\operatorname{sgn}(g_{\theta}(X_{i}))=\left|\psi_{i}\right|(1-2\mathbb{I}\left[\operatorname{sgn}(g_{\theta}(X_{i}))\neq\operatorname{sgn}(\psi_{i})\right]$). Then we can use surrogate losses for the zero-one loss,
$L(\theta)=\mathbb{E}[\left|\psi\right|\ell(g_{\theta}(X),\operatorname{sgn}(\psi))]$
Although many functional forms for $\ell(\cdot)$ are Fisher-consistent, the logistic (cross-entropy) loss will be particularly relevant: $l(g,s)=2\log(1+\exp(g))-(s+1)g$.
Example 8 (Treatment parity, continued (weighted classification reduction)).
The cost-sensitive reduction for a vector of Lagrange multipliers can be deduced by applying the weighted classification reduction to the Lagrangian:
$L(\beta)=\mathbb{E}\left[|\tilde{\psi}^{\lambda}|\ell\left(f_{\beta}(X),\operatorname{sgn}(\tilde{\psi}^{\lambda})\right)\right],\qquad\text{ where }\tilde{\psi}^{\lambda}=\psi+\frac{\lambda_{A}}{p_{A}}(p_{1\mid 1}-p_{1\mid 0})-\sum_{a\in\mathcal{A}}\lambda_{a}.$
where $p_{a}:=\hat{P}(A=a)$ and $\lambda_{a}:=\lambda_{(a,+)}-\lambda_{(a,-)}$, effectively replacing two non-negative Lagrange multipliers by a single multiplier, which can be either positive or negative.
Example 9 (Responder-conditional treatment parity, continued).
The Lagrangian is $L(\beta)=\mathbb{E}\left[|\tilde{\psi}^{\lambda}|\ell\left(f_{\beta}(X),\operatorname{sgn}(\tilde{\psi}^{\lambda})\right)\right]$ with weights:
$\tilde{\psi}^{\lambda}=\psi+\frac{\lambda_{A}}{p_{A}}\frac{(p_{1\mid 1}-p_{1\mid 0})(\mu_{1}-\mu_{0})}{\mathbb{E}_{n}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]}-\sum_{a\in\mathcal{A}}\lambda_{a}.$
where $p_{a}:=\hat{P}(A=a)$ and $\lambda_{a}:=\lambda_{(a,+)}-\lambda_{(a,-)}$.
C.3 Proofs
Proof of Proposition 1.
$\mathbb{E}[c(\pi,T(\pi),Y(\pi))]$ $=\textstyle\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[\mathbb{I}\left[T(r)=t\right]c_{rt}(Y(t(r)))\mid R=r,X]]$
$\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X){P}(T=t\mid R=r,X)\mathbb{E}[c_{rt}(Y(t(r)))\mid R=r,X]]$
$\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X){P}(T=t\mid R=r,X)\mathbb{E}[c_{rt}(Y)\mid T=t,X]]$
where the last line follows by the conditional exclusion restriction (Assumption 2) and consistency (Assumption 1).
Proof of Proposition 7.
$P(T(\pi)=1\mid A=a,Y(1(\pi))>Y(0(\pi)))$
$=\frac{P(T(\pi)=1,Y(1(r))>Y(0(r))\mid A=a)}{P(Y(1(\pi))>Y(0(\pi))\mid A=a)}$ by Bayes’ rule
$=\frac{P(T(\pi)=1,Y(1)>Y(0)\mid A=a)}{P(Y(1)>Y(0)\mid A=a)}$ by Assumption 2
$=\frac{\sum_{r}\mathbb{E}[\mathbb{E}[\pi_{r}(X)\mathbb{I}\left[T(r)=1\right]\mathbb{I}\left[Y(1)>Y(0)\right]\mid A=a,X]])}{P(Y(1)>Y(0)\mid A=a)}$ by iter. exp
$\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]}$ by Proposition 1
Appendix D Proofs
D.1 Proofs for generalization under unconstrained policies
Proposition 8 (Policy value generalization).
Assume the nuisance models $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0},e_{r}(X)]^{\top},\eta\in\mathcal{F}_{\eta}$ are consistent and well-specified with finite VC-dimension $v_{\eta}$ over the product function class $H$. We provide a proof for the general case, including doubly-robust estimators, which applies to the statement of Proposition 8 by taking $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0}].$
Let $\Pi=\{\mathbb{I}\{\mathbb{E}[L(\lambda,X,A;\eta)\mid X]<0\}\colon\lambda\in\mathbb{R};\eta\in\mathcal{F}_{\eta}\}.$
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi L(\lambda,X,A)]-\mathbb{E}[\pi L(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
The generalization bound allows deducing risk bounds on the out-of-sample value:
Corollary 4.
$\mathbb{E}[L(\hat{\lambda},X,A)_{+}]\leq\mathbb{E}[L(\lambda^{*},X,A)_{+}]+O_{p}(n^{-\frac{1}{2}})$
Proof of Proposition 8.
We study a general Lagrangian, which takes as input pseudo-outcomes $\psi^{t\mid r}(O;\eta),\psi^{y\mid t}(O;\eta),\psi^{1\mid 0,\Delta A}$ where each satisfies that
$\mathbb{E}[\psi^{t\mid r}(O;\eta)\mid X,A]$ $=p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A)$
$\mathbb{E}[\psi^{y\mid t}(O;\eta)\mid X,A]$ $=\tau(X,A)$
$\mathbb{E}[\psi^{1\mid 0,\Delta A}\mid X]$ $=p_{1\mid 0}(X,a)-p_{1\mid 0}(X,b)$
We make high-level stability assumptions on pseudooutcomes $\psi$ relative to the nuisance functions $\eta$ (these are satisfied by standard estimators that we will consider):
Assumption 11.
$\psi^{t\mid r},\psi^{y\mid t},\psi^{1\mid 0,\Delta A}$ respectively are Lipschitz contractions with respect to $\eta$ and bounded
We study a generalized Lagrangian of an optimization problem that took these pseudooutcome estimates as inputs:
$L(\lambda,X,A;\eta)=\psi_{t\mid r}(O;\eta)\left\{\psi_{y\mid t}(O;\eta)+\frac{\lambda}{p(A)}(\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right])\right\}+\lambda(\psi^{1\mid 0,\Delta A}(O;\eta))$
We will show that
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi L(\lambda,X,A)]-\mathbb{E}[\pi L(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
which, by applying the generalization bound twice gives that
$\mathbb{E}_{n}[\pi L(\lambda,X,A)]=\mathbb{E}[\pi L(\lambda,X,A)])+O_{p}(n^{-\frac{1}{2}})$
Write Lagrangian as
$\max_{\pi}\min_{\lambda}=\min_{\lambda}\max_{\pi}=\min_{\lambda}\mathbb{E}[L(O,\lambda;\eta)_{+}]$
Suppose the Rademacher complexity of $\eta_{k}$ is given by $\mathcal{R}(H_{k}),$ so that [Bartlett and Mendelson, 2002, Thm. 12] gives that the Rademacher complexity of the product nuisance class $H$ is therefore $\sum_{k}\mathcal{R}(H_{k}).$ The main result follows by applying vector-valued extensions of Lipschitz contraction of Rademacher complexity given in Maurer [2016]. Suppose that $\psi^{t\mid r},\psi^{y\mid t},\psi^{1\mid 0,\Delta A}$ are Lipschitz with constants $C^{L}_{t\mid r},C^{L}_{y\mid t},C^{L}_{1\mid 0,\Delta A}$.
We establish VC-properties of
$\mathcal{F}_{L_{1}}(O_{1:n})=\left\{(g_{\eta}(O_{1}),g_{\eta}(O_{i}),\dots g_{\eta}(O_{n}))\colon\eta\in\mathcal{F}_{\eta}\right\},\text{ where}g_{\eta}(O)=\psi_{t\mid r}(O;\eta)\psi_{y\mid t}(O;\eta)$
$\mathcal{F}_{L_{2}}(O_{1:n})=\left\{(h_{\eta}(O_{1}),h_{\eta}(O_{i}),\dots h_{\eta}(O_{n}))\colon\eta\in\mathcal{F}_{\eta}\right\},\text{ where}h_{\eta}(O)=\psi_{t\mid r}(O;\eta)\frac{\lambda}{p(A)}(\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right])$
$\mathcal{F}_{L_{3}}(O_{1:n})=\left\{(m_{\eta}(O_{1}),m_{\eta}(O_{i}),\dots m_{\eta}(O_{n}))\colon\eta\in\mathcal{F}_{\eta}\right\},\text{ where}m_{\eta}(O)=\lambda(\psi^{1\mid 0,\Delta A}(O;\eta))$
and the function class for the truncated Lagrangian,
$\mathcal{F}_{L_{+}}=\left\{\{(g_{\eta}(O_{i})+h_{\eta}(O_{i})+m_{\eta}(O_{i}))_{+}\}_{1:n}\colon g\in\mathcal{F}_{L_{1}}(O_{1:n}),h\in\mathcal{F}_{L_{2}}(O_{1:n}),m\in\mathcal{F}_{L_{3}}(O_{1:n}),\eta\in\mathcal{F}_{\eta}\right\}$
[Maurer, 2016, Corollary 4] (and discussion of product function classes) gives the following: Let $\mathcal{X}$ be any set, $\left(x_{1},\ldots,x_{n}\right)\in\mathcal{X}^{n}$, let $F$ be a class of functions $f:\mathcal{X}\rightarrow\ell_{2}$ and let $h_{i}:\ell_{2}\rightarrow\mathbb{R}$ have Lipschitz norm $L$. Then
$\mathbb{E}\sup_{\eta\in H}\sum_{i}\epsilon_{i}\psi_{i}\left(\eta\left(O_{i}\right)\right)\leq\sqrt{2}L\mathbb{E}\sup_{\eta\in H}\sum_{i,k}\epsilon_{ik}\eta\left(O_{i}\right)\leq\sqrt{2}L\sum_{k}\mathbb{E}\sup_{\eta_{k}\in H_{k}}\sum_{i}\epsilon_{i}\eta_{k}\left(O_{i}\right)$ (5)
where $\epsilon_{ik}$ is an independent doubly indexed Rademacher sequence and $f_{k}\left(x_{i}\right)$ is the $k$-th component of $f\left(x_{i}\right)$.
Applying Equation 5 to each of the component classes $\mathcal{F}_{L_{1}}(O_{1:n}),\mathcal{F}_{L_{2}}(O_{1:n}),\mathcal{F}_{L_{3}}(O_{1:n}),$ and Lipschitz contraction [Bartlett and Mendelson, 2002, Thm. 12.4] of the positive part function $\mathcal{F}_{L_{+}}$, we obtain the bound
$\sup_{\lambda,\eta}\left|\mathbb{E}_{n}[L(O,\lambda;\eta)_{+}]-\mathbb{E}[L(O,\lambda;\eta)_{+}]\right|\leq\sqrt{2}(C^{L}_{t\mid r}C^{L}_{y\mid t}+C^{L}_{t\mid r}B_{p_{a}}B+BC^{L}_{1\mid 0,\Delta A})\sum_{k}\mathcal{R}(H_{k})$
Proposition 9 (Threshold solutions).
Define
$L(\lambda,X,A)=(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\left\{\tau(X,A)+\frac{\lambda}{p(A)}(\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right])\right\}+\lambda(p_{1\mid 0}(X,a)-p_{1\mid 0}(X,b))$
$\lambda^{*}\in\arg\max_{\lambda}\;\mathbb{E}[L(\lambda,X,A)_{+}],\;\;\pi^{*}(x,u)=\mathbb{I}\{L(\lambda^{*},X,u)<0\}$
If instead $d(x)$ is a function of covariates $x$ only,
$\lambda^{*}\in\arg\max_{\lambda}\;\mathbb{E}[\mathbb{E}[L(\lambda,X,A)\mid X]_{+}],\;\;\pi^{*}(x)=\mathbb{I}\{\mathbb{E}[L(\lambda^{*},X,A)\mid X]<0\}$
Proof of Proposition 9.
The characterization follows by strong duality in infinite-dimensional linear programming [Shapiro, 2001]. Strict feasibility can be satisfied by, e.g. solving LABEL:eqn-feas-constr to set ranges for $\epsilon$. ∎
D.2 Proofs for robust characterization
Proof of Proposition 6.
$V(\pi)=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[c_{rt}(Y(t))\mathbb{I}\left[T(r)=t\right]\mid R=r,X]]$
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[c_{rt}(Y(t))\mid R=r,X]P(T(r)=t\mid R=r,X)]$ unconf.
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[c_{rt}(Y(t))\mid X]P(T(r)=t\mid R=r,X)]$ Assumption 2 (ER) (6)
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\mathbb{E}\left[c_{rt}(Y(t))\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}\mid X\right]P(T(r)=t\mid R=r,X)\right]$ unconf.
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\mathbb{E}\left[c_{rt}(Y(t))\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}+\left(1-\frac{T}{p_{t}(X)}\right)\mu_{t}(X)\mid X\right]p_{t\mid r}(X)\right\}\right]$ control variate
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\left\{c_{y}{rt}(Y(t))\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}+\left(1-\frac{T}{p_{t}(X)}\right)\mu_{t}(X)\right\}p_{t\mid r}(X)\right\}\right]$ (LOTE)
where $p_{t}(X)=P(T=t\mid X)$ (marginally over $R$ in the observational data) and (LOTE) is an abbreviation for the law of total expectation. ∎
Proof of Lemma 1.
$\overline{V}_{no}(\pi)$ $\coloneqq\underset{q_{tr}(X)\in\mathcal{U}}{\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mu_{t}(X)q_{tr}(X)\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]\right\}$
$=\underset{q_{tr}(X)\in\mathcal{U}}{\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[Y\mid T=t,X]q_{tr}(X)\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]\right\}$
Note the objective function can be reparametrized under a surjection of $q_{t\mid r}(X)$ to its marginalization, i.e. marginal expectation over a $\{T=t\}$ partition (equivalently $\{T=t,A=a\}$ partition for a fairness-constrained setting).
Define
$\beta_{t\mid r}(a)\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t,A=a],\beta_{t\mid r}\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t]$
Therefore we may reparametrize $\overline{V}_{no}(\pi)$ as an optimization over constant coefficients (bounded by B):
$={\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\{c_{t}\beta_{t\mid r}\}\pi_{r}(X)\mathbb{E}[Y\mid T=t,X]\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]\colon\underline{B}\leq c_{1}\leq\overline{B},c_{0}=1-c_{1}\right\}$
$={\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\{c_{t}\beta_{t\mid r}\}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]\colon\underline{B}\leq c_{1}\leq\overline{B},c_{0}=1-c_{1}\right\}$ (LOTE)
$=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[c^{*}_{rt}\beta_{t\mid r}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]$
[…]
Proof of proposition 5.
$\min_{\pi}\;$ $\mathbb{E}[c(\pi,T(\pi),Y(\pi))\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]+\mathbb{E}[c(\pi,T(\pi),Y(\pi))\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]$ (7)
$\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}\mid A=b]$ (8)
$\qquad\;\;+\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}\mid A=b]\leq\epsilon,\forall q_{r1}\in\mathcal{U}$ (9)
Define
$g_{r}(x,u)=(\mu_{r1}(x,u)-\mu_{r0}(x,u))$
then we can rewrite this further and apply the standard epigraph transformation:
$\min_{t,\pi}\;$ $t$
$t-\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{r1}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}],\forall q_{r1}\in\mathcal{U}$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)q_{r1}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)q_{r1}(x,b))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,\forall q_{r1}\in\mathcal{U}$
Project the uncertainty set onto the direct product of uncertainty sets:
$\min_{t,\pi}\;$ $t$
$t-\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{r1}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}],\forall q_{r1}\in\mathcal{U_{1}}$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)q_{r1}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)q_{r1}(x,b))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,\forall q_{r1}\in\mathcal{U_{2}}$
Clearly robust feasibility of the resource parity constraint over the interval is obtained by the highest/lowest bounds for groups $a,b$, respectively:
$\min_{t,\pi}\;$ $t$
$t-\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{r1}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}],\forall q_{r1}\in\mathcal{U_{1}}$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
We define
$\delta_{r1}(x,u)=\frac{2(q_{r1}(x,u)-\underline{B}_{r}(x,u))}{\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)}-(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)),$
then
$\{\underline{B}_{r}(x,u)\leq q_{r1}(x,u)\leq\overline{B}_{r}(x,u)\}\implies\{\lVert\delta_{r1}(x,u)\rVert_{\infty}\leq 1\}$
and
$q_{r1}(x,u)=\underline{B}_{r}(x,u)+\frac{1}{2}(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u))(\delta_{r1}(x,u)+1).$
For brevity we denote $\Delta B=(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u))$, so
$\min_{t,\pi}\;t$
$t+\min_{\begin{subarray}{c}\lVert\delta_{r1}(x,u)\rVert_{\infty}\leq 1\\ r\in\{0,1\},u\in\{a,b\}\end{subarray}}\left\{-\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}\frac{1}{2}\Delta B(x,u)\delta_{r1}(x,u)dx\right\}$
$\qquad\qquad-c_{1}(\pi)\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}]$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,$
where
$c_{1}(\pi)=\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}(\underline{B}_{r}(x,u)+\frac{1}{2}(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)))dx$
This is equivalent to:
$\min_{t,\pi}\;$ $t$
$t+\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}-\left|g_{r}(x,u)\pi_{r}(x,u)f(x,u)\right|\frac{1}{2}\Delta B(x,u)dx-c_{1}(\pi)\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}]$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
Undoing the epigraph transformation, we obtain:
$\min_{\pi}\;$ $V_{ov}(\pi)+\mathbb{E}[\mu_{0}]+c_{1}(\pi)+\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\left|g_{r}(x,u)\pi_{r}(x,u)f(x,u)\right|\frac{1}{2}\Delta B(x,u)dx$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
and simplifying the absolute value:
$\min_{\pi}\;$ $V_{ov}(\pi)+\mathbb{E}[\mu_{0}]+c_{1}(\pi)+\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\left|g_{r}(x,u)\pi_{r}(x,u)f(x,u)\right|\frac{1}{2}\Delta B(x,u)dx$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
D.3 Proofs for general fairness-constrained policy optimization algorithm and analysis
We begin by summarizing some notation that will simplify some statements. Define, for observation tuples $O\sim(X,A,R,T,Y),$ the value estimate $v(Q;\eta)$ given some pseudo-outcome $\psi(O;\eta)$ dependent on observation information and nuisance functions $\eta$. (We often suppress notation of $\eta$ for brevity). We let estimators sub/super-scripted by $1$ denote estimators from the first dataset.
$v_{(\cdot)}(O;Q,\eta)$ $=\mathbb{E}_{\pi\sim Q}[v_{(\cdot)}(O;\pi,\eta)]$
$v_{(\cdot)}(Q)$ $=\mathbb{E}[v_{(\cdot)}(Q)]$
$\hat{V}_{1}^{(\cdot)}(Q)$ $=\mathbb{E}_{n_{1}}[v_{(\cdot)}(Q)]$
$g_{j}(O;Q)$ $=\mathbb{E}_{\pi\sim Q}[g_{j}(O;\pi)\mid O,\mathcal{E}_{j}]$
$h_{j}(Q)$ $=\mathbb{E}[g_{j}(O;Q)\mid\mathcal{E}_{j}]$
$\hat{h}^{1}_{j}(Q)$ $=\mathbb{E}_{n_{1}}[g_{j}(O;Q)\mid\mathcal{E}_{j}]$
D.3.1 Preliminaries: results from other works used without proof
Theorem 5 (Theorem 3, [Agarwal et al., 2018] (saddle point generalization bound for LABEL:alg-metaalg1) ).
Let $\xi:=\max_{h}\|{M}\hat{{\mu}}(h)-\hat{{c}}\|_{\infty}$. Suppose Assumption 7 holds for $C^{\prime}\geq 2C+2+\sqrt{\ln(4/\delta)/2}$, where $\delta>0$. Let $Q^{\star}$ minimize $V(Q)$ subject to ${M}{h}(Q)\leq{c}$. Then Algorithm 1 with $\nu\propto n^{-\alpha},B\propto n^{\alpha}$ and $\omega\propto\xi^{-2}n^{-2\alpha}$ terminates in $O\left(\xi^{2}n^{4\alpha}\ln|\mathcal{K}|\right)$ iterations and returns $\hat{Q}$. If $np_{j}^{\star}\geq 8\log(2/\delta)$ for all $j$, then with probability at least $1-(|\mathcal{J}|+1)\delta$ then for all $k$, $\hat{Q}$ satisfies:
$V(\hat{Q})\leq V\left(Q^{\star}\right)+\widetilde{O}\left(n^{-\alpha}\right)$
$\gamma_{k}(\widehat{Q})\leq d_{k}+\frac{1+2\nu}{B}+\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\widetilde{O}\left(\left(np_{j}^{\star}\right)^{-\alpha}\right)$
The proof of [Agarwal et al., 2018, Thm. 3] is modular in invoking Rademacher complexity bounds on the objective function and constraint moments, so that invoking standard Rademacher complexity bounds for off-policy evaluation/learning [Athey and Wager, 2021, Swaminathan and Joachims, 2015] yields the above statement for $V(\pi)$ (and analogously, randomized policies by [Bartlett and Mendelson, 2002, Thm. 12.2] giving stability for convex hulls of policy classes).
More specifically, we use standard local Rademacher complexity bounds.
Definition 1 (Local Rademacher Complexity).
The local Rademacher complexity for a generic $f\in\mathcal{F}$ is:
$\mathcal{R}(r,\mathcal{F})=\mathbb{E}_{\epsilon,X_{1:n}}\left[\sup_{f\in\mathcal{F}:\|f\|_{2}\leq r}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f\left(X_{i}\right)\right]$
The following is a generic concentration inequality for local Rademacher complexity over some radius $r$; see Wainwright [2019] for more background.
Lemma 2 (Lemma 5 of [Chernozhukov et al., 2019]/Lemma 4, Foster and Syrgkanis [2019]).
Consider any $Q^{*}\in\mathcal{Q}$. Assume that $v(\pi)$ is L-Lipschitz in its first argument with respect to the $\ell_{2}$ norm and let:
$Z_{n}(r)=\sup_{Q\in\mathcal{Q}}\{\left|\mathbb{E}_{n}[\hat{v}(Q)-\hat{v}(Q^{*})]-\mathbb{E}[v(Q)-v(Q^{*})]\right|\colon\mathbb{E}[(v(Q)-v(Q^{*}))^{2}]^{\frac{1}{2}}\leq r\}$
Then for some constant $C_{3}$:
$Z_{n}(r)\leq C_{3}\left(\mathcal{R}\left(r,\mathcal{Q}-Q^{*}\right)+r\sqrt{\frac{\log(1/\delta)}{n}}+\frac{\log(1/\delta)}{n}\right)$
Lemma 3 (Concentration of conditional moments ([Agarwal et al., 2018, Woodworth et al., 2017]).
For any $j\in\mathcal{J}$, with probability at least $1-\delta$, for all $Q$,
$\left|\widehat{h}_{j}(Q)-h_{j}(Q)\right|\leq 2\mathcal{R}_{n_{j}}(\mathcal{Q})+\frac{2}{\sqrt{n_{j}}}+\sqrt{\frac{\ln(2/\delta)}{2n_{j}}}$
If $np_{j}^{\star}\geq 8\log(2/\delta)$, then with probability at least $1-\delta$, for all $Q$,
$\left|\widehat{h}_{j}(Q)-h_{j}(Q)\right|\leq 2\mathcal{R}_{np_{j}^{\star}/2}(\mathcal{Q})+2\sqrt{\frac{2}{np_{j}^{\star}}}+\sqrt{\frac{\ln(4/\delta)}{np_{j}^{\star}}}$
Lemma 4 (Orthogonality (analogous to [Chernozhukov et al., 2019] (Lemma 8), others)).
Suppose the nuisance estimates satisfy a mean-squared-error bound
$\max_{l}\{\mathbb{E}[(\hat{\eta}_{l}-\eta_{l})^{2}]\}_{l\in[L]}\coloneqq\chi_{n}^{2}$
Then w.p. $1-\delta$ over the randomness of the policy sample,
$V(Q_{0})-V(\hat{Q})\leq O(R_{n,\delta}+\chi^{2}_{n})$
D.4 Adapted lemmas
In this subsection we collect results similar to those that have appeared previously, but that require substantial additional argumentation in our specific saddle point setting.
The next lemma establishes the variance of small-regret policies with estimated vs. true nuisances is close, up to nuisance estimation error.
Lemma 5 (Feasible vs. oracle nuisances in low-variance regret slices (Chernozhukov et al. [2019], Lemma 9) ).
Suppose that the mean squared error of the nuisance estimates is upper bounded w.p. $1-\delta$ by $\chi_{n,\delta}^{2}$ and suppose $\chi_{n,\delta}^{2}\leq\epsilon_{n}$. Then:
$V_{2}^{0}=\sup_{Q,Q^{\prime}\in\mathcal{Q}_{*}\left(\epsilon_{n}+2\chi_{n,\delta}^{2}\right)}\operatorname{Var}\left(v_{DR}^{0}(x;Q)-v_{DR}^{0}\left(x;Q^{\prime}\right)\right)$
Then $V_{2}\leq V_{2}^{0}+O\left(\chi_{n,\delta}\right)$.
D.5 Proof of Theorem 3
Proof of Theorem 3.
We first study the meta-algorithm with “oracle" nuisance functions $\eta=\eta_{0}$. For brevity below we notationally suppress the dependence of $v$ on observation $O$.
Define
$\Pi_{2}\left(\epsilon_{n}\right)$ =$=\left\{\pi\in{\Pi}:\mathbb{E}_{n_{1}}[v(\pi;\eta_{0})-v(\hat{Q}_{1};\eta_{0})]\leq\epsilon_{n},\mathbb{E}_{n_{1}}\left[g_{j}(O;\pi,\eta_{0})-g_{j}\left(O;\hat{Q}_{1},\eta_{0}\right)\mid\mathcal{E}_{j}\right]\leq\epsilon_{n},j\in\hat{\mathcal{I}}_{1}\right\}$
$\mathcal{Q}_{2}\left(\epsilon_{n}\right)$ =$=\left\{Q\in\Delta\left({\Pi_{2}}(\epsilon_{n})\right)\right\}$
$\mathcal{Q}^{*}\left(\epsilon_{n}\right)$ =$=\left\{Q\in\Delta({\Pi}):\mathbb{E}[(v(Q;\eta_{0})-v(Q^{*};\eta_{0})]\leq\epsilon_{n},\mathbb{E}[g_{j}(O;Q,\eta_{0})\mid\mathcal{E}_{j}]-\mathbb{E}[g_{j}(O;Q^{*},\eta_{0})\mid\mathcal{E}_{j}]\leq\epsilon_{n}\right\}$
In the following, we suppress notational dependence on $\eta_{0}.$
Note that $\hat{Q}_{1}\in\mathcal{Q}_{2}\left(\epsilon_{n}\right).$
Step 1: First we argue that w.p. $1-\delta/6,$ $Q^{*}\in\mathcal{Q}_{2}$.
Invoking Theorem 5 on the output of the first stage of the algorithm, yields that with probability $1-\frac{\delta}{6}$ over the randomness in $\mathcal{D}_{1},$ by choice of $\epsilon_{n}$ = $\bar{O}(n^{-\alpha}))$,
$V(\hat{Q}_{1})$ $\leq V(Q^{*})+\epsilon_{n}/2$
$\gamma_{k}(\hat{Q}_{1})$ $\leq d_{k}+\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\widetilde{O}\left((np_{j}^{*})^{-\alpha}\right)\leq d_{k}+\epsilon_{n}/2\quad\text{ for all }k$
Further, by Lemma 2,
$\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[(v(Q)-v(Q^{*}))]-\mathbb{E}[(v(Q)-v(Q^{*}))]\right|\leq\epsilon_{n}/2$
$\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[(g(O;Q)-g(O;Q^{*}))]-\mathbb{E}[(g(O;Q)-g(O;Q^{*}))]\right|\leq\epsilon_{n}/2$
Therefore, with high probability on the good event, $Q^{*}\in\mathcal{Q}_{2}.$
Step 2: Again invoking Theorem 5, this time on the output of the second stage of the algorithm with function space $\Pi_{2}$ (hence implicitly $\mathcal{Q}_{2}$), and conditioning on the “good event" that $Q^{*}\in\mathcal{Q}_{2}$, we obtain the bound that with probability $\geq 1-\delta/3$ over the randomness of the second sample $\mathcal{D}_{2},$
$V(\hat{Q}_{2})$ $\leq V(Q^{*})+\epsilon_{n}/2$
$\gamma_{k}(\hat{Q}_{2})$ $\leq\gamma_{k}({Q}^{*})+\epsilon_{n}/2$
Step 3: empirical small-regret slices relate to population small-regret slices, and variance bounds
We show that if $Q\in\mathcal{Q}_{2},$ then with high probability $Q\in\mathcal{Q}_{2}^{0}$ (defined on small population value- and constraint-regret slices relative to $\hat{Q}_{1}$ rather than small empirical regret slices)
$\mathcal{Q}_{2}^{0}=\{Q\in\operatorname{conv}(\Pi)\colon\left|V(Q)-V(\hat{Q}_{1})\right|\leq\epsilon_{n}/2,\mathbb{E}[g_{j}(O;Q)-g_{j}(O;\hat{Q}_{1}))\mid\mathcal{E}_{j}]\leq\epsilon_{n},\forall j\}$
so that w.h.p. $\mathcal{Q}_{2}\subseteq\mathcal{Q}_{2}^{0}.$
Note that for $Q\in\mathcal{Q}$, w.h.p. $61-\delta/$ 6 over the first sample, we have that
$\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n}[v(Q)-v(\hat{Q}_{1})]-\mathbb{E}[v(Q)-v(\hat{Q}_{1})]\right|\leq 2\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n}[v(Q)]-\mathbb{E}[v(Q)]\right|\leq\epsilon,$
$\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[g_{j}(O;Q)-g_{j}(O;\hat{Q}_{1})\mid\mathcal{E}_{j}]-\mathbb{E}[g_{j}(O;Q)-g_{j}(O;\hat{Q}_{1})\mid\mathcal{E}_{j}]\right|$
$\qquad\qquad\leq 2\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[g_{j}(O;Q)\mid\mathcal{E}_{j}]-\mathbb{E}[g_{j}(O;Q)\mid\mathcal{E}_{j}]\right|\leq\epsilon,\forall j$
The second bound follows from [Bartlett and Mendelson, 2002, Theorem 12.2] (equivalence of Rademacher complexity over convex hull of the policy class) and linearity of the policy value and constraint estimators in $\pi,$ and hence $Q$.
On the other hand since $Q_{1}$ achieves low policy regret, the triangle inequality implies that we can contain the true policy by increasing the error radius. That is, for all $Q\in\mathcal{Q}_{2},$ with high probability $\geq 1-\delta/3$:
$\left|\mathbb{E}[(v(Q)-v(Q^{*}))]\right|\leq\left|\mathbb{E}[(v(Q)-v(\hat{Q}_{1}))]\right|+\left|\mathbb{E}[(v(\hat{Q}_{1})-v(Q^{*}))]\right|\leq\epsilon_{n}$
$\left|\mathbb{E}[g_{j}(O;Q)-g_{j}(O;Q^{*})\mid\mathcal{E}_{j}]\right|\leq\left|\mathbb{E}[g_{j}(O;Q)-g_{j}(O;\hat{Q}_{1})\mid\mathcal{E}_{j}]\right|+\left|\mathbb{E}[g_{j}(O;\hat{Q}_{1})-g_{j}(O;Q^{*})\mid\mathcal{E}_{j}]\right|\leq\epsilon_{n}$
Define the space of distributions over policies that achieve value and constraint regret in the population of at most $\epsilon_{n}:$
$\mathcal{Q}_{*}(\epsilon_{n})=\{Q\in\mathcal{Q}\colon V(Q)-V(Q^{*})\leq\epsilon_{n},\;\;\mathbb{E}[g_{j}(O;Q)-g_{j}(O;Q^{*})\mid\mathcal{E}_{j}]\leq\epsilon_{n},\forall j\},$
so that on that high-probability event,
$\mathcal{Q}_{2}^{0}(\epsilon_{n})\subseteq\mathcal{Q}_{*}(\epsilon_{n}).$ (10)
Then on that event with probability $\geq 1-\delta/3$,
$r_{2}^{2}$ $=\sup_{Q\in\mathcal{Q}_{2}}\mathbb{E}[(v(Q)-v(Q^{*}))^{2}]\leq\sup_{Q\in\mathcal{Q}^{*}(\epsilon_{n})}\mathbb{E}[(v(Q)-v(Q^{*}))^{2}]$
$=\sup_{Q\in\mathcal{Q}^{*}(\epsilon_{n})}\operatorname{Var}(v(Q)-v(Q^{*}))+\mathbb{E}[(v(Q)-v(Q^{*}))]^{2}$
$\leq\sup_{Q\in\mathcal{Q}^{*}(\epsilon_{n})}\operatorname{Var}(v(Q)-v(Q^{*}))+\epsilon_{n}^{2}$
Therefore:
$r_{2}\leq\sqrt{\sup_{Q\in\mathcal{Q}_{*}\left(\epsilon_{n}\right)}\operatorname{Var}\left(v(Q)-v(Q^{*})\right)}+2\epsilon_{n}=\sqrt{V_{2}}+2\epsilon_{n}$
Combining this with the local Rademacher complexity bound, we obtain that:
$\mathbb{E}[v(\hat{Q}_{2})-v(Q^{*})]=O\left(\kappa\left(\sqrt{V_{2}}+2\epsilon_{n},\mathcal{Q}_{*}\left(\epsilon_{n}\right)\right)+\sqrt{\frac{V_{2}\log(3/\delta)}{n}}\right)$
These same arguments apply for the variance of the constraints
$V_{2}^{j}=\sup\;\{\operatorname{Var}\left(g_{j}(O;Q)-g_{j}\left(O;Q^{\prime}\right)\right)\colon{Q,Q^{\prime}\in\mathcal{Q}_{*}(\tilde{\epsilon}_{n})}\}$
D.6 Proofs of auxiliary/adapted lemmas
Proof of Lemma 5.
The proof is analogous to that of [Chernozhukov et al., 2019, Lemma 9] except for the step of establishing that $\pi_{*}\in\mathcal{Q}^{0}_{\epsilon_{n}+O(\chi_{n,\delta}^{2})}$: in our case we must establish relationships between saddlepoints under estimated vs. true nuisances. We show an analogous version below.
Define the saddle points to the following problems (with estimated vs. true nuisances):
$(Q^{*}_{0,0},\lambda^{*}_{0,0})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[v_{DR}(Q;\eta_{0})]+\lambda^{\top}(\gamma_{DR}(Q;\eta_{0})-d)\coloneqq L(Q,\lambda;\eta_{0},\eta_{0})\coloneqq L(Q,\lambda),$
$(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[v_{DR}(Q;\eta)]+\lambda^{\top}(\gamma_{DR}(Q;\eta_{0})-d),$
$(Q^{*},\lambda^{*})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[v_{DR}(Q;\eta)]+\lambda^{\top}(\gamma_{DR}(Q;\eta)-d).$
We have that:
$\mathbb{E}[v_{DR}(Q^{*})]\leq L(Q^{*},\lambda^{*};\eta,\eta)+\nu$
$\leq L(Q^{*},\lambda^{*};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$
$\leq L(Q^{*},\lambda^{*};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$ by Lemma 4
$\leq L(Q^{*},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$ by saddlepoint prop.
$\leq L(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\left|L(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})-L(Q^{*},\lambda^{*}_{\eta,0};\eta,\eta_{0})\right|+\nu+\chi_{n,\delta}^{2}$ triangle ineq.
$\leq L(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\epsilon_{n}+\nu+\chi_{n,\delta}^{2}$ assuming $\text{assuming }\epsilon_{n}\geq\chi_{n,\delta}^{2}$
$\leq\mathbb{E}[v_{DR}(Q^{*}_{\eta,0};\eta)]+\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}$ apx. complementary slackness
$\leq\mathbb{E}[v_{DR}(Q^{*}_{0,0};\eta)]+\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}$ suboptimality
Hence
$\mathbb{E}[v_{DR}(Q^{*};\eta)]-\mathbb{E}[v_{DR}(Q^{*}_{0,0};\eta)]\leq\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}.$
We generally assume that the saddlepoint suboptimality $\nu$ is of lower order than $\epsilon_{n}$ (since it is under our computational control).
Applying Lemma 4 gives;
$V(Q^{*})-V(Q^{*}_{0,0})\leq\epsilon_{n}+2\nu+2\chi_{n,\delta}^{2}.$
Define policy classes with respect to small-population regret slices (with a nuisance-estimation enlarged radius):
$\mathcal{Q}^{0}(\epsilon)=\{Q\in\Delta(\Pi)\colon V(Q^{*}_{0})-V(Q)\leq\epsilon,\gamma(Q^{*}_{0})-\gamma(Q)\leq\epsilon\}$
Then we have that
$V_{2}^{obj}\leq\sup_{Q\in\mathcal{Q}^{0}(\epsilon_{n})}\operatorname{Var}(v_{DR}(O;\pi)-v_{DR}(O;\pi^{*})),$
where we have shown that $\pi^{*}\in\mathcal{Q}^{0}(\epsilon+2\nu+2\chi_{n,\delta}^{2}).$
Following the rest of the argumentation in [Chernozhukov et al., 2019, Lemma 9] from here onwards gives the result, i.e. studying the case of estimated nuisances with our Lemma 5 and Lemma 4. ∎
Appendix E Additional case study
E.1 Oregon Health Insurance Study
The Oregon Health Insurance Study [Finkelstein et al., 2012] is an important study on the causal effect of expanding public health insurance on healthcare utilization, outcomes, and other outcomes. It is based on a randomized controlled trial made possible by resource limitations, which enabled the use of a randomized lottery to expand Medicaid eligibility for low-income uninsured adults. Outcomes of interest included health care utilization, financial hardship, health, and labor market outcomes and political participation.
Because the Oregon Health Insurance Study expanded access to enroll in Medicaid, a social safety net program, the effective treatment policy is in the space of encouragement to enroll in insurance (via access to Medicaid) rather than direct enrollment. This encouragement structure is shared by many other interventions in social services that may invest in nudges to individuals to enroll, tailored assistance, outreach, etc., but typically do not automatically enroll or automatically initiate transfers. Indeed this so-called administrative burden of requiring eligible individuals to undergo a costly enrollment process, rather than automatically enrolling all eligible individuals, is a common policy design lever in social safety net programs. Therefore we expect many beneficial interventions in this consequential domain to have this encouragement structure.
We preprocess the data by partially running the Stata replication file, obtaining a processed data file as input, and then selecting a subset of covariates that could be relevant for personalization. These covariates include household information that affected stratified lottery probabilities, socioeconomic demographics, medical status and other health information.
In the notation of our framework, the setup of the optimal/fair encouragement policy design question is as follows:
  • $X$: covariates (baseline household information, socioeconomic demographics, health information)
  • $A$: race (non-white/white), or gender (female/male)
    These protected attributes were binarized.
  • $R$: encouragement: lottery status of expanded eligibility (i.e. invitation to enroll when individual was previously ineligible to enroll)
  • $T$: whether the individual is enrolled in insurance ever
    Note that for $R=1$ this can be either Medicaid or private insurance while for $R=0$ this is still well-defined as this can be private insurance.
  • $Y$: number of doctor visits
    This outcome was used as a measure of healthcare utilization. Overall, the study found statistically significant effects on healthcare utilization. An implicit assumption is that increased healthcare utilization leads to better health outcomes.
We subsetted the data to include complete cases only (i.e. without missing covariates). We learned propensity and treatment propensity models via logistic regression for each group, and used gradient-boosted regression for the outcome model. We first include results for regression adjustment identification. One potential concern is the continued use of the healthcare utilization variable as an outcome measure. From a methodological angle, it displays heterogeneity in treatment effects. From the substantive angle, healthcare utilization remains a proxy outcome measure for other health measures, and interpreting increases in healthcare utilization as beneficial is justified primarily by assuming that individuals were constrained by the costs of uninsured healthcare previously, so that increases in healthcare utilization reflect that access to insurance increases in access to care.
Figure 7: Distribution of lift in treatment probabilities $p_{1\mid 1,a}-p_{1\mid 0,a}=P(T=1\mid R=1,A=a,X)-P(T=1\mid R=0,A=a,X)$, and plot of $p_{1\mid 1,a}-p_{1\mid 0,a}$ vs. $\tau.$
Figure 8: Policy value $V(\pi^{\lambda})$, treatment value $\mathbb{E}[T(\pi^{\lambda})\mid A=a],$ for $A$ = race, gender.
In Figure 8 we plot descriptive statistics. We include histograms of the treatment responsivity lifts $p_{1\mid 1a}(x,a)-p_{1\mid 0a}(x,a).$ We see some differences in distributions of responsivity by gender and race. We then regress treatment responsivity on the outcome-model estimate of $\tau$. We find substantially more heterogeneity in treatment responsivity by race than by gender: whites are substantially more likely to take up insurance when made eligible, conditional on the same expected treatment effect heterogeneity in increase in healthcare utilization. (This is broadly consistent with health policy discussions regarding mistrust of the healthcare system).
Next we consider imposing treatment parity constraints on an unconstrained optimal policy (defined on these estimates). In Figure 8 we display the objective value, and $\mathbb{E}[T(\pi)\mid A=a]$, for gender and race, respectively, enumerated over values of the constraint. We use costs of $2$ for the number of doctors visits and $1$ for enrollment in Medicaid (so $\mathbb{E}[T(\pi)\mid A=a]$ is on the scale of probability of enrollment). These costs were chosen arbitrarily. Finding optimal policies that improve disparities in group-conditional access can be done with relatively little impact to the overall objective value. These group-conditional access disparities can be reduced from $4$ percentage points ($0.04$) for gender and about $6$ percentage points $(0.06)$ for race at a cost of $0.01$ or $0.02$ in objective value (twice the number of doctors’ visits). On the other hand, relative improvements/compromises in access value for the “advantaged group" show different tradeoffs. Plotting the tradeoff curve for race shows that, consistent with the large differences in treatment responsivity we see for whites, improving access for blacks. Looking at this disparity curve given $\lambda$ however, we can also see that small values of $\lambda$ can have relatively large improvements in access for blacks before these improvements saturate, and larger $\lambda$ values lead to smaller increases in access for blacks vs. larger decreases in access for whites.
E.2 Additional Discussion, PSA-DMF case study
First, before describing the analysis, we acknowledge important data issues (such as those that commonly arise from the criminal justice system [Bao et al., 2021]), in addition to particularities of this data set, so that this analysis should be viewed as exploratory.
Our analysis proceeds conditional on the non-detained population. This could make sense in a setting where decision-making frameworks for supervised release are unlikely to change judicial decisions to detain or release: our results apply to marginal defendants. Covariate levels (including PSA scores) were discretized for privacy. Moreover, the recorded final supervision decision does not include intensity, but different intensities are recommended in the data, which we collapse into a single level. The PSA-DMF is an algorithmic recommendation so here we are appealing to overlap in treatment recommendations, but using parametric extrapolation in responsivity. So, we are assuming randomness in treatment assignment that arises either from quasi-random assignment to judges or noise/variability in judicial decisions. We strongly appeal to this interpretation of randomness in treatment assignment in the conceptualization of a causal effect of treatment with supervised release. Other accounts and conceptualizations of judicial decision-making could instead argue that judicial decisions such as conditional release are by their very nature discretionary, and do not admit valid counterfactuals. We instead appeal to a hypothetical randomized experiment (if unethical) where individuals could conceivably be randomized into supervised release or not. Finally, unconfoundedness is likely untrue, but sensitivity analysis could address this in ways quite similar to those studied previously in the literature [Kallus et al., 2019b, Kallus and Zhou, 2021b, 2018].